### v7

# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

Abstract
We tackle the Optimal Experiment Design Problem, which consists of choosing experiments to run or observations to select from a finite set to estimate the parameters of a system. The objective is to maximize some measure of information gained about the system from the observations, leading to a convex integer optimization problem. We leverage Boscia.jl, a recent algorithmic framework, which is based on a nonlinear branch-and-bound algorithm with node relaxations solved to approximate optimality using Frank-Wolfe algorithms. One particular advantage of the method is its efficient utilization of the polytope formed by the original constraints which is preserved by the method, unlike alternative methods relying on epigraph-based formulations. We assess the method against both generic and specialized convex mixed-integer approaches. Computational results highlight the performance of the proposed method, especially on large and challenging instances.
1 Introduction
The Optimal Experiment Design Problem (OEDP) arises in statistical estimation and empirical studies in many applications areas from Engineering to Chemistry. For OEDP, we assume we have a matrix $A$ consisting of the rows $\mathbf{v}_{1},\allowbreak\dots,\mathbf{v}_{m}\in{\mathbb{R}}^{n}$ where each row represents an experiment. The ultimate aim is to fit a regression model:
$\min_{\bm{\theta}\in{\mathbb{R}}^{m}}\|A\bm{\theta}-\mathbf{y}\|,$ (1)
where $\mathbf{y}$ encodes the responses of the experiments and $\bm{\theta}$ are the parameters to be estimated. The set of parameters with size $n$ is assumed to be (significantly) smaller than the number of distinct experiments $m$. Furthermore, we assume that $A$ has full column rank, i.e. the vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{m}$ span ${\mathbb{R}}^{n}$.
The problem is running all experiments, potentially even multiple times to account for errors, is often not realistic because of time and cost constraints. Thus, OEDP deals with finding a subset of size $N$ of the experiments providing the “most information” about the experiment space (pukelsheim2006optimal; de1995d). In general, the number of allowed experiments $N$ is in the interval $[n,m]$ in order to allow a solution to the regression model in Equation˜1.
In Section˜2, we investigate the necessary conditions for a function to be a valid and useful information measure. Every information function leads to a different criterion. In this paper, we focus on two popular criteria, namely the A-criterion and D-criterion, see Section˜2.1.
In general, OEDP leads to a Mixed-Integer Non-Linear Problem (MINLP). There has been a lot of development in the last years in solving MINLP (kronqvist2019review). Nevertheless, the capabilities of current MINLP solvers are far away from their linear counterparts, the Mixed-Integer Problem (MIP) solvers (bestuzheva2021scip), especially concerning the magnitude of the problems that can be handled. Therefore, instead of solving the actual MINLP formulation, a continuous version of OEDP, called the Limit Problem, is often solved and the integer solution is created from the continuous solution by rounding schemes and other heuristics (pukelsheim2006optimal). This does not necessarily lead to optimal solutions, though, and the procedure is not always applicable to a given continuous solution either because it requires a certain sparsity of the continuous solution.
The goal of this paper is to compare the performance of different MINLP approaches for OEDP problems. A special focus is put on the newly proposed framework Boscia (hendrych2023convex) which can solve larger instances and significantly outperforms the other examined approaches. The new method leverages Frank-Wolfe algorithms and a formulation different from the other approaches. In Section˜3, we establish the convergence of the Frank-Wolfe algorithm on the continuous relaxations of the considered OEDP problems. The different solution methods are detailed in Section˜4 and evaluation of the computational experiments can be found in Section˜5.
1.1 Related work
As mentioned, one established method of solution is the reduction to a simpler problem by removing the integer constraints and employing heuristics to generate an integer solution from the continuous solution. Recently, there have been more publications tackling the MINLP formulation of the Optimal Experiment Design Problem. These, however, concentrate on specific information measures, in particular the A-criterion, see (nikolov2022proportional; ahipacsaouglu2015first), and the D-criterion, (welch1982branch; ponte2023branch; li2022d; xu2024submodular; ponte2023computing; fampa2022maximum; chen2024generalized; ponte2024convex). The most general solution approach known to the authors was introduced in ahipacsaouglu2021branch. It considers OEDP under matrix means which, in particular, includes the A-criterion and D-criterion. While the matrix means covers many information measures of interest, it still yields a restricting class of information functions. For example, the G-criterion and V-criterion are not included in this class of functions (de1995d). The newly proposed framework Boscia only requires the information measures to be either $L$-smooth, i.e. the gradient is Lipschitz continuous, or generalized self-concordant, thereby covering a larger group of information functions. In addition, Boscia does not suppose any prior knowledge about the structure of the problem, being thus more flexible in terms of problem formulations. On the other hand, it is highly customizable, giving the user the ability to exploit the properties of their problems to speed up the solving process. An in-depth and unified theory for the Optimal Experiment Design Problem can be found in pukelsheim2006optimal.
1.2 Contribution
Our contribution can be summarized as follows
Unified view on experiment design formulations.
First, we propose a unified view of multiple experiment design formulations as the optimization of a nonlinear (not necessarily Lipschitz-smooth) information function over a truncated scaled probability simplex intersected with the set of integers. Unlike most other formulations that replace the nonlinear objective with nonlinear and/or conic constraints, we preserve the original structure of the problem. Additionally, we can easily handle special cases of OEDP without any reformulations and additional constraints on OEDP since we do not suppose a specific problem structure, unlike the approach in ahipacsaouglu2021branch.
Convergence analysis for BPCG.
Second, we establish convergence guarantees for Frank-Wolfe algorithms on continuous relaxations of these problems by exploiting smoothness or self-concordance, and sharpness. This is necessary in order to solve the various experimental design problems with Boscia.
Superior solution via Boscia framework.
Finally, we use the recently proposed Boscia (hendrych2023convex) framework to solve the integer optimization problem with a Frank-Wolfe method for the node relaxations of the Branch-and-Bound tree and show the effectiveness of the method on instances generated with various degrees of correlation between the parameters.
1.3 Notation
In the following let $\lambda_{i}(A)$ denote the $i$-th eigenvalue of matrix $A$; we assume that these are sorted in increasing order. Moreover, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the minimum and maximum eigenvalue of $A$, respectively. We define and denote the condition number of a positive semi-definite matrix $A$ as $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$. Further, let $\log\det(\cdot)$ be the log-determinant of a positive definite matrix. Given matrices $A$ and $B$ of same dimensions, $A\circ B$ denotes their Hadamard product. Given a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes the diagonal matrix with $\mathbf{x}$ on its diagonal and let $\operatorname{Diag}(A)$ denote the diagonal of the matrix $A$. The cones of positive definite and positive semi-definite matrices in ${\mathbb{R}}^{n\times n}$ will be denoted by $\mathbb{S}_{++}^{n}$ and $\mathbb{S}_{+}^{n}$, respectively. We will refer to them as PD and PSD cones. The set of symmetric $n\times n$ matrices will be denoted by $\mathbb{S}^{n}$. Let ${\mathbb{N}}_{0}$ denote the set of positive integers including 0 and for $m\in{\mathbb{N}}$ let $m]=\{1,2,\dots,m\}$. Lastly, we denote matrices with capital letters, e.g. $X$, vectors with bold small letters, e.g. $\mathbf{x}$, and simple small letters are scalars, e.g. $\lambda$.
2 The optimal experiment design problem
As stated in the introduction, OEDP aims to pick the subset of predefined size $N$ yielding the most information about the system. So first, we have to answer the question of how to quantify information. To that end, we introduce the information matrix
$X(\mathbf{x})=\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}=A^{\intercal}\operatorname{diag}(\mathbf{x})A$ (2)
where $x_{i}\in{\mathbb{N}}_{0}$ denotes the number of times experiment $i$ is to be performed. We call $\mathbf{x}\in{\mathbb{N}}_{0}^{m}$ a design. Throughout this paper, we will use both ways of expressing $X(\mathbf{x})$ but will favor the second representation. The inverse of the information matrix is the dispersion matrix
$D(\mathbf{x})=\left(\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}\right)^{-1}.$
It is a measure of the variance of the experiment parameters (ahipacsaouglu2015first). Our goal is to obtain the “most” information, i.e. we want to maximize some measure of the information matrix. This is equivalent to minimizing over the dispersion matrix (pukelsheim2006optimal), thereby minimizing the variance of the parameters.
[…]
Remark.
Experiments can be run only once or be allowed to run multiple times to account for measurement errors. In the latter case, we will suppose non-trivial upper (and lower bounds) on the number of times a given experiment can be run. The sum of the upper bounds significantly exceeds $N$.
If non-trivial lower bounds $\mathbf{l}$ are present, their sum may not exceed $N$ otherwise there is no solution respecting the time and cost constraints. A special case of non-trivial lower bounds is obtained if $n$ linearly independent experiments have non-zero lower bounds. These experiments can be summarized in the matrix $C=A^{\intercal}\operatorname{diag}(\mathbf{l})A$. Notice that $C$ is positive definite. The information matrix then becomes
$X_{C}(\mathbf{x})=C+A^{\intercal}\operatorname{diag}(\mathbf{x}-\mathbf{l})A.$ (3)
We will refer to the resulting OEDP as the Fusion Experiment Design Problem or short as the Fusion Problem. The OEDP with $\mathbf{l}=0$ will be referred to as Optimal Problem. These are the two types of problems we are focusing on in this paper.
Let us now answer the question posed earlier: How can we measure information? We need a function $\phi$ receiving a positive definite matrix [^2] As previously stated, the information matrix of the optimal solution $X(\mathbf{x}^{*})$ has to have full column rank. Thus it suffices to define $\phi$ on the PD cone. as input and returning a number, that is $\phi:\mathbb{S}_{++}^{n}\rightarrow{\mathbb{R}}$. We will lose information by compressing a matrix to a single number. Hence, the suitable choice of $\phi$ depends on the underlying problem. Nevertheless, there are some properties that any $\phi$ has to satisfy to qualify as an information measure.
Definition 1 (Information Function (pukelsheim2006optimal)).
An information function $\phi$ on $\mathbb{S}_{++}^{n}$ is a function $\phi:\mathbb{S}_{++}^{n}\rightarrow{\mathbb{R}}$ that is positively homogeneous, concave, nonnegative, non-constant, upper semi-continuous and respects the Loewner ordering.
Respecting Loewner ordering
Let $B,D\in\mathbb{S}_{+}^{n}$. We say $D\succcurlyeq B$ if and only if $D-B\succcurlyeq 0$ which is equivalent to $D-B\in\mathbb{S}_{+}^{n}$. The order thus imposed on $\mathbb{S}_{+}^{n}$ is called the Loewner Ordering. A map $\phi$ respects the Loewner ordering iff $C\succcurlyeq B\Rightarrow\phi(C)\geq\phi(B)\;\forall C,B\in\mathbb{S}_{+}^{n}$. In simple terms, carrying out more experiments should not result in information loss.
Concavity
The condition for concavity is
$\phi(\gamma B+(1-\gamma)C)\geq\gamma\phi(B)+(1-\gamma)\phi(C)\quad\forall\gamma\in[0,1]\;\forall B,C\in\mathbb{S}_{++}^{n}.$
The intuition behind this is that we should not be able to achieve the best result by performing two sets of experiments and interpolating the result. There should be a design unifying the two sets achieving a better result.
Positive homogeneity
For $\gamma>0$, we have $\phi(\gamma B)=\gamma\phi(B)$, i.e. scaling has no influence on the overall information value. It has added advantage that later in the optimization, we may forgo any factors since they do not interfere with the ordering created by $\phi$.
Nonnegativity
By convention, we have $\phi(B)\geq 0$ for all $B\in\mathbb{S}_{+}^{n}$.
Upper semi-continuity
The upper level sets $\left\{M\in\mathbb{S}_{++}\,|\,\phi(M)\geq\lambda\right\}$ are closed for all $\lambda\in{\mathbb{R}}$. This is to ensure "good" behavior at the boundary. The functions considered in this paper are all continuous.
The most frequently used information functions arise from matrix means $\phi_{p}$ (pukelsheim2006optimal, ahipacsaouglu2015first). The matrix means are defined as follows:
Definition 2 (Matrix mean).
Let $C\in\mathbb{S}_{++}^{n}$ and let $\lambda(C)$ denote the eigenvalues of $C$. The matrix mean $\phi_{p}$ of $C$ is defined as
$\phi_{p}(C)=\begin{cases}\max\lambda(C),&\text{for }p=\infty\\ \left(\frac{1}{n}\operatorname{Tr}(C^{p})\right)^{\frac{1}{p}},&\text{for }p\neq 0,\pm\infty\\ \det(C)^{\frac{1}{n}},&\text{for }p=0\\ \min\lambda(C),&\text{for }p=-\infty.\end{cases}$ (4)
If $C$ is semi-definite with $\operatorname{rank}(C)<n$, the matrix mean is defined as:
$\phi_{p}(C)=\begin{cases}\max\lambda(C),&\text{for }p=\infty\\ \left(\frac{1}{n}\operatorname{Tr}(C^{p})\right)^{\frac{1}{p}},&\text{for }p=(0,\infty)\\ 0,&\text{for }p=[0,-\infty].\end{cases}$ (5)
Note that the function $\phi_{p}$ satisfies the requirements of Definition˜1 only for $p\leq 1$ (pukelsheim2006optimal). The different values of $p$ lead to different so-called criteria. The Optimal Experiment Design Problem using a matrix means function is then defined as
$\max_{\mathbf{x}}\,$ $\log(\phi_{p}(X(\mathbf{x})))$ (OEDP)
s.t. $\sum_{i=1}^{m}x_{i}=N$
$\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}$
$\mathbf{x}\in{\mathbb{N}}_{0}^{m},$
where $\mathbf{u}$ and $\mathbf{l}$ denote the upper and lower bounds, respectively.
Remark.
Using the logarithm formulation, one can recover Fenchel duality results (sagnol2010optimal, Chapter 3). It can also by beneficial from a information theory point of view. Note, however, that the $\log$ does not necessarily preserve concavity.
The resulting problem (OEDP) is an integer optimization problem which, depending on the information function $\phi_{p}$, can be $\mathcal{NP}$-hard. The two information measures we will focus on lead to $\mathcal{NP}$-hard problems (welch1982branch; nikolov2022proportional; li2025strong; li2022d). Thus, these OEDP are hard to solve. Often, therefore, one solves the so-called Limit Problem obtained by letting the number of allowed experiments $N$ go to infinity. The resulting optimization problem is continuous.
$\max\limits_{\mathbf{w}}\,$ $\log(\phi_{p}(X(\mathbf{w})))$ (OEDP-Limit)
s.t. $\sum_{i=1}^{m}w_{i}=1$
$\mathbf{w}\in[0,1]^{m}$
The variable vector $\mathbf{w}$ can be interpreted as a probability distribution. Under suitable assumptions, one can generate a finite experiment design from the solution of the limit problem, see ahipacsaouglu2015first, pukelsheim2006optimal. One assumption is that the support of the solution of the Limit Problem (OEDP-Limit) is smaller than the allowed number of experiments $N$. This is often not the case (pukelsheim2006optimal). If an integer solution can be obtained, it is feasible but not necessarily optimal. Optimality can be achieved in some special cases, see pukelsheim2006optimal. However, these have very specific requirements that are often not achievable.
2.1 The A-, D- and GTI-optimal experiment design problem
For this paper, we focus on two particular criteria arising from the matrix means. The two most commonly used criteria are the D-optimality and A-optimality criterion, $p=0$ and p=1p=-1, respectively.
Many methods for solving MINLP expect minimization formulations. Hence, we will reformulate the problems as minimization problems by flipping the sign of the objective.
D-Criterion
Choosing $p=0$ in Definition˜2, yields for the objective of (OEDP)
$\max\limits_{\mathbf{x}\in{\mathbb{N}}_{0}^{m}}\,\log\det\left((X(\mathbf{x}))^{\frac{1}{n}}\right)$
Note that $\log(\det(X)^{\frac{1}{n}})=\frac{1}{n}\log\det X$ and thus we can state the D-Optimal Experiment Design Problem as:
$\begin{split}\min\limits_{\mathbf{x}}\,&-\log(\det(X(\mathbf{x})))\\ \text{s.t. }\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{Z}}\end{split}$ (D-Opt)
Observe that it is equivalent to minimize $\log\det(X(\mathbf{x})^{-1})=\log\det D(\mathbf{x})$, so the determinant of the dispersion matrix. This is also called the generalized variance of the parameter $\bm{\theta}$ (pukelsheim2006optimal). Thus, a maximal value of $\det X$ corresponds to a maximal volume of standard ellipsoidal confidence region of $\bm{\theta}$ (ponte2023branch). Additionally, the D-criterion is invariant under reparameterization (pukelsheim2006optimal).
A-Criterion
For parameters with a physical interpretation, the A-optimality criterion is a good choice (pukelsheim2006optimal) as it amounts to minimizing the average of the variances of $\bm{\theta}$. Geometrically, the A-criterion amounts to minimizing the diagonal of the bounding box of the ellipsoidal confidence region (sagnol2010optimal). The resulting optimization problem can be stated in two different ways, keeping the $\log$ and using the $\log$ rules and the positive homogeneity of the information functions.
$\begin{split}\min\limits_{\mathbf{x}}\,&\log\left(\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (logA-Opt)
$\begin{split}\min\limits_{\mathbf{x}}\,&\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (A-Opt)
We test both formulation within the newly proposed framework. Both formulations can be generalized for a real number $p>0$:
$\begin{split}\min\limits_{\mathbf{x}}\,&\log\left(\operatorname{Tr}\left((X(\mathbf{x}))^{-p}\right)\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (logGTI-Opt)
$\begin{split}\min\limits_{\mathbf{x}}\,&\operatorname{Tr}\left((X(\mathbf{x}))^{-p}\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (GTI-Opt)
We call the corresponding problem the Generalized-Trace-Inverse Optimal Experiment Design Problem, short GTI-Optimal Problem.
Remark.
We conjecture that OEDP under the GTI-criterion is $\mathcal{NP}$-hard for any $p>0$.
Note that the objectives of (D-Opt), (A-Opt), (logA-Opt), (logGTI-Opt) and (GTI-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank.
Remark.
The Fusion Problems are created by replacing $X(\mathbf{x})$ with $X_{C}(\mathbf{x})$ in (D-Opt), (A-Opt), (logA-Opt), (logGTI-Opt) and (GTI-Opt), respectively.
For convenience in proofs in the next section, we define
$\mathcal{P}:=[\mathbf{l},\mathbf{u}]\cap\left\{\mathbf{x}\in{\mathbb{R}}^{m}_{\geq 0}\,\middle|\,\sum x_{i}=N\right\}$ (Convex Hull)
as the feasible region of the OEDPs without the integer constraints, i.e. the convex hull of all feasible integer points. Thus, the feasible region with integer constraints will be noted by $\mathcal{P}\cap{\mathbb{Z}}$.
We denote the domain of the objective functions by
$D:=\left\{\mathbf{x}\in{\mathbb{R}}^{m}_{\geq 0}\,\middle|\,X(\mathbf{x})\in\mathbb{S}_{++}^{n}\right\}$ (Domain)
Throughout the rest of this paper, we denote by
$f(X)=-\log\det(X)$
$g(X)=\operatorname{Tr}\left(X^{-p}\right)$
$k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$
where $X\in\mathbb{S}_{++}^{n}$ and $p>0$. With abusive of notation, we use the same identifiers for the objectives of the optimization problems.
$f(\mathbf{x})=-\log\det(X(\mathbf{x}))$
$g(\mathbf{x})=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$
$k(\mathbf{x})=\log\left(\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)\right)$
3 Properties of the problems and convergence guarantees
To ensure that the newly proposed framework in hendrych2023convex is applicable to the previously introduced problems, we have to show that the Frank-Wolfe algorithms used as the subproblem solvers converge on the continuous subproblems. Frank-Wolfe normally requires the objective function to be convex and $L$-smooth. The convexity part is trivial, for proofs see Section˜A.2. The $L$-smoothness property only holds for the Fusion Problems on the whole feasible region. For the Optimal Problems, we will show local $L$-smoothness in Section˜3.1 and an alternative property guaranteeing convergence of Frank-Wolfe, namely (generalized) self-concordance, in Section˜3.2.
Lastly, we establish sharpness or the Hölder error bound condition of the objective functions, therefore yielding improved convergence rates for Frank-Wolfe methods compared to the standard $\mathcal{O}(1/\epsilon)$ rate where $\epsilon$ is the additive error in the primal gap, see Section˜3.3.
3.1 Lipschitz smoothness
The convergence analysis for Frank-Wolfe algorithms relies on the fact that the objective function values do not grow arbitrarily large. The property guaranteeing this is Lipschitz smoothness, further called $L$-smoothness.
Definition 3.
A function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is called $L$-smooth if its gradient $\nabla f$ is Lipschitz continuous. That is, there exists a finite $L\in{\mathbb{R}}_{>0}$ with
$\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq L\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$
Since the functions of interest are in $\mathcal{C}^{3}$, we can use an alternative condition on $L$-smoothness:
$LI\succcurlyeq\nabla_{xx}f(\mathbf{x}),$
i.e. the Lipschitz constant $L$ is an upper bound on the maximum eigenvalue of the Hessian.
In Section˜A.3, the reader can find $L$-smoothness proofs for $f(X)$, $g(X)$ and $k(X)$ over
$\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\geq\lambda_{\min}(X)\}.$
The lemmas are not immediately useful for us, in general we cannot bound the minimum eigenvalue of the information matrix. In fact, we will see that we have $L$-smoothness on the whole convex feasible region $\mathcal{P}$ only for the Fusion Problems. Due to the positive definiteness of the matrix $C$, the information matrix in the Fusion case, $X_{C}(\mathbf{x})$, is always bounded away from the boundary of the PD cone and we do find a lower bound on the minimum eigenvalue of $X_{C}(\mathbf{x})$. For Optimal Problems, $X(\mathbf{x})$ can be singular for some $\mathbf{x}\in\mathcal{P}$, i.e. the minimum eigenvalue of $X(\mathbf{x})$ cannot be bounded away from 0 on $\mathcal{P}$. However, we can show $L$-smoothness in a local region for the Optimal Problems.
We start with the $L$-smoothness proofs for the objectives of the Fusion Problems.
Theorem 1.
The functions $f_{F}(\mathbf{x}):=-\log\det\left(X_{C}(\mathbf{x})\right)$ and $g_{F}(\mathbf{x}):=\operatorname{Tr}\left(\left(X_{C}(\mathbf{x})\right)^{-p}\right)$ for $p\in{\mathbb{R}}_{>0}$ are $L$-smooth on $x\in{\mathbb{R}}_{\geq 0}$ with Lipschitz constants
$L_{f_{F}}=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}}{\lambda_{\min}(C)^{2}}$
and
$L_{g_{F}}=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}},$
respectively.
Proof.
For better readability and simplicity, we ignore the dependency of $X$ on $\mathbf{x}$ for now. As stated above, we can prove $L$-smoothness by showing that the largest eigenvalue of the Hessians is bounded for any $\mathbf{x}$ in the domain. We have
$\nabla^{2}f_{F}(\mathbf{x})$ $=(AA^{\intercal})\circ(AX_{C}^{-2}A^{\intercal})$
and
$\nabla^{2}g_{F}(\mathbf{x})$ $=p(p+1)\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right).$
Observe that the Hessian expressions are very similar. Hence, we will only present the proof for $g_{F}$. The proof of $f_{F}$ follows the same argumentation. By horn1990hadamard it is known that the maximum eigenvalue of a Hadamard product between two positive semi-definite matrices $N$ and $M$, $\lambda_{\max}(N\circ M)$, is upper bounded by the maximum diagonal entry of $N$ and the largest eigenvalue of $M$, so
$\lambda_{\max}(N\circ M)$ $\leq\max_{i}N_{ii}\lambda_{\max}(M).$ (8)
In our case, $N=AA^{\intercal}$ and $M=AX_{C}^{-p-2}A^{\intercal}$. By Corollary˜2 in Section˜A.1, we have that $X_{C}^{-p-2}$ is positive definite. Thus, both factors are semi-definite and we may use the provided bound.
$\lambda_{\max}\left(\nabla^{2}g_{F}(\mathbf{x})\right)$ $\leq\max\limits_{i}(AA^{\intercal})_{ii}\lambda_{\max}(AX_{C}^{-p-2}A^{\intercal})$
Note that the first factor is simply
$\max\limits_{i}(AA^{\intercal})_{ii}$ $=\max\limits_{i}\mathbf{v}_{i}^{\intercal}\mathbf{v}_{i}.$ (9)
For the second factor, we can exploit positive semi-definiteness with the spectral norm.
$\lambda_{\max}\left(AX_{C}^{-p-2}A^{\intercal}\right)$ $=\left\|AX_{C}^{-p-2}A^{\intercal}\right\|_{2}$
$\leq\left\|A\right\|_{2}^{2}\left\|X_{C}^{-p-2}\right\|_{2}$
By Lemma˜6, we have $\lambda(A^{p})=\lambda(A)^{p}$.
$=\left\|A\right\|_{2}^{2}\lambda_{\max}(X_{C}^{-1})^{2+p}$
$=\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X_{C})^{2+p}}$
By the definition of $X$ and by Weyl’s inequality on the eigenvalues of sums of positive semi-definite matrices weyl1912asymptotische, we have that
$\lambda_{\min}(X_{C})$ $\geq\lambda_{\min}(C)+\lambda_{\min}(A^{\intercal}\operatorname{diag}(x)A)\geq\lambda_{\min}(C).$
Hence,
$\lambda_{\max}\left(AX_{C}^{-p-2}A^{\intercal}\right)$ $\leq\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}.$ (10)
Combining (8) with (9) and (10) yields
$\lambda_{\max}\left(\nabla^{2}g_{F}(\mathbf{x})\right)$ $\leq p(p+1)\left(\max\limits_{i}v_{i}^{\intercal}v_{i}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}.$ (11)
For the $\log$ variant of the A-criterion, we can also bound the maximum eigenvalue of the hessian $\nabla^{2}k_{F}$.
Theorem 2.
The function $k_{F}(\mathbf{x}):=\log\operatorname{Tr}\left(\left(X_{C}(\mathbf{x})\right)^{-p}\right)$ is $L$-smooth on $\mathcal{P}$ with Lipschitz constant
$L_{k_{F}}=\frac{a_{C}^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}$
where $a_{C}$ is the upper bound on the maximum eigenvalue of $X_{C}(\mathbf{x})$, that is
$a_{C}=\lambda_{\max}(C)+\max_{i\in[m]}u_{i}\max_{j\in[m]}\left\|v_{j}\right\|_{2}^{2}$
with $u_{i}$ denoting the upper bound of variable $x_{i}$.
Proof.
The Hessian of $k_{F}$ is given by
$\nabla^{2}k_{F}$ $=p\frac{(p+1)\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)\operatorname{Tr}\left(X_{C}^{-p}\right)-p\operatorname{diag}\left(AX_{C}^{-p-1}A^{\intercal}\right)\operatorname{diag}\left(AX_{C}^{-p-1}A^{\intercal}\right)^{\intercal}}{\operatorname{Tr}\left(X_{C}^{-p}\right)^{2}}.$
As before, we omit the dependency on $\mathbf{x}$ for better readability. Note that both parts of the nominator are symmetric matrices. Thus, we can use the upper bound on the eigenvalues of sums of symmetric matrices in Lemma˜15. Further, note that for any $\mathbf{z}\in{\mathbb{R}}^{n}$ the matrix $\mathbf{z}\mathbf{z}^{\intercal}$ is PSD and its maximum eigenvalue is $\left\langle\mathbf{z},\mathbf{z}\right\rangle$.
$\lambda_{\max}\left(\nabla^{2}k_{F}\right)$ $\leq\frac{p(p+1)}{\operatorname{Tr}\left(X_{C}^{-p}\right)}\lambda_{\max}\left(\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)\right)$
An upper bound on the maximum eigenvalue of $\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)$ is already provided by Theorem˜1. Observe that
$\operatorname{Tr}\left(X_{C}^{-p}\right)\geq n\lambda_{\min}\left(X_{C}^{-p}\right)=\frac{n}{\lambda_{\max}(X_{C})^{p}}\geq\frac{n}{a_{C}}.$
Assembling everything yields the desired bound. ∎
As stated before, the proofs of Theorems˜1 and 2 work because we can bound the minimum eigenvalue of the Fusion information matrix $X_{C}$ away from zero thanks to the regular matrix $C$. This is not the case of the information matrices $X(\mathbf{x})$ of the Optimal Problems as they can be singular for some $x\in\mathcal{P}$. Therefore, the objectives of the Optimal Problems are not $L$-smooth over $\mathcal{P}$. We can show, however, that the objectives are $L$-smooth on a restricted area of the feasible region.
Remember that $D$ denotes the domain of the objectives, see (Domain). Since the Frank-Wolfe algorithms make monotonic progress, it is helpful to show that given an initial point $\mathbf{x}_{0}$ in the domain $D$ the minimum eigenvalue is bounded in the region
$\mathcal{L}_{0}=\left\{\mathbf{x}\in D\cap\sum\limits_{i=1}^{m}x_{i}=N\,\middle|\,(*)(\mathbf{x})\leq(*)(\mathbf{x}_{0})\right\}$
where ()(*) is a placeholder for $f(\mathbf{x})$, $g(\mathbf{x})$ and $k(\mathbf{x})$, respectively. Using the bound provided by merikoski1997bounds and the fact that the objective value is decreasing, we can bound the minimum eigenvalue by quantities depending on the experiment matrix $A$ and the start point $\mathbf{x}_{0}$. Let further $X_{0}=X(\mathbf{x}_{0})$.
Theorem 3.
The functions $f(\mathbf{x}):=-\log\det\left(X(\mathbf{x})\right)$, $g(\mathbf{x}):=\operatorname{Tr}\left(\left(X(\mathbf{x})\right)^{-p}\right)$ and $k(\mathbf{x}):=\operatorname{Tr}\left(\left(X(\mathbf{x})\right)^{-p}\right)$ for $p>0$ are $L$-smooth locally on their respective $\mathcal{L}_{0}$ for a given initial point $x_{0}$ with
$L_{f}=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}\left(mN\max_{i}\left\|v_{i}\right\|_{2}^{2}\right)^{n-1}}{\left(n-1\right)^{n-1}\det X_{0}^{2}}$
$L_{g}=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
and
$L_{k}=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
where $a=\max_{i\in[m]}u_{i}\max_{j\in[m]}\left\|v_{j}\right\|_{2}^{2}$ is a upper bound on the maximum eigenvalue of $X(\mathbf{x})$.
Proof.
As seen in the previous proofs, we have
$L_{f}$ $=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2}}$
$L_{g}$ $=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2+p}}.$
and
$L_{k}$ $=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2+p}}$
To bound the Lipschitz constant, we need lower bound the minimum eigenvalue of $X(\mathbf{x})$. We start with $L_{f}$ and find by using the bound provided by merikoski1997bounds
$\lambda_{\min}(X)$ $\geq\left(\frac{n-1}{\operatorname{Tr}(X)}\right)^{n-1}\det X.$
The trace can be upper bounded by $\operatorname{Tr}(X)\leq mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}$ and by knowing that $\log\det(X(\mathbf{x}))\leq-\log\det(X_{0})$, we find
$\lambda_{\min}(X)$ $\geq\left(\frac{n-1}{mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}}\right)^{n-1}\det X_{0}.$
Thus,
$L_{f}$ $=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}\left(mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}\right)^{n-1}}{\left(n-1\right)^{n-1}\det X_{0}^{2}}.$
To lowerbound the minimum eigenvalue of $X(\mathbf{x})$ w.r.t. the objectives $g(\mathbf{x})$ and $k(\mathbf{x})$, we first need
$\operatorname{Tr}(X^{-p-2})$ $=\sum\limits_{i=1}^{n}\frac{1}{\lambda_{i}(X)^{p+2}}=\sum\limits_{i=1}^{n}\frac{1}{\lambda_{i}(X)^{2}}\frac{1}{\lambda_{i}(X)^{p}}$
$\leq\frac{1}{\lambda_{\min}(X)^{2}}\operatorname{Tr}(X^{-p}).$
Note that the $\log$ is preserving ordering. Hence, we can use the fact that $\operatorname{Tr}(X^{-p})\leq\operatorname{Tr}(X_{0}^{-p})$ and we find
$\leq\frac{1}{\lambda_{\min}(X)^{2}}\operatorname{Tr}(X_{0}^{-p}).$
Exploiting positive definiteness and the spectral norm, we have
$\lambda_{\min}(X)^{p+2}$ $=\lambda_{\min}\left(X^{\frac{p+2}{2}}\right)^{2}=\frac{1}{\lambda_{\max}\left(X^{-\frac{p+2}{2}}\right)^{2}}$
$=\frac{1}{\left\|X^{-\frac{p+2}{2}}\right\|_{2}^{2}}\geq\frac{1}{\left\|X^{-\frac{p+2}{2}}\right\|_{F}^{2}}=\frac{1}{\operatorname{Tr}\left(X^{-p-2}\right)}$
$\lambda_{\min}(X)^{p+2}$ $\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}(X^{-p})}\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}\left(X_{0}^{-p}\right)}\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}\left(X_{0}^{-1}\right)^{p}}.$
Ultimately, we find
$\lambda_{\min}(X)$ $\geq\frac{1}{\operatorname{Tr}\left(X_{0}^{-1}\right)}$
and the smooth constants are
$L_{g}$ $=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
and
$L_{k}$ $=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}.$
Thereby, we have local smoothness constants for the objectives $f$, $g$ and $k$. ∎
To summarize, we have shown $L$-smoothness of the objectives of the Fusion Problems which guarantees convergence of Frank-Wolfe. For the Optimal Problems, we have shown local $L$-smoothness on a restricted area of the feasible region. This, however, does not suffice to guarantee convergence of Frank-Wolfe. In the next section, we will show that the objectives satisfy an alternative condition that ensures convergence of Frank-Wolfe.
3.2 Generalized self-concordance
By carderera2021simple, the Frank-Wolfe algorithm also converges (with similar convergence rates to the $L$-smooth case) if the objective is generalized self-concordant.
Definition 4 (Generalized Self-Concordance sun2019generalized).
A three-times differentiable, convex function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is $(M_{f},\nu)$-generalized self-concordant with order $\nu>0$ and constant $M_{f}\geq 0$, if for all $x\in\text{dom}(f)$ and $\mathbf{u},\mathbf{v}\in{\mathbb{R}}^{n}$, we have
$\left|\left\langle\nabla^{3}f(\mathbf{x})[\mathbf{u}]\mathbf{v},\mathbf{v}\right\rangle\right|\leq M_{f}\|\mathbf{u}\|^{2}_{x}\;\|\mathbf{v}\|_{x}^{\nu-2}\;\|\mathbf{v}\|_{2}^{3-\nu}$
where
$\|\mathbf{w}\|_{x}=\left\langle\nabla^{2}f(\mathbf{x})\mathbf{w},\mathbf{w}\right\rangle\text{ and }\nabla^{3}f(\mathbf{x})[\mathbf{u}]=\lim\limits_{\gamma\rightarrow 0}\gamma^{-1}\left(\nabla^{2}f(\mathbf{x}+\gamma\mathbf{u})-\nabla^{2}f(\mathbf{x})\right).$
Self-concordance is a special case of generalized self-concordance where $\nu=3$ and $\mathbf{u}=\mathbf{v}$. This yields the condition
$\left|\left\langle\nabla^{3}f(\mathbf{x})[\mathbf{u}]\mathbf{u},\mathbf{u}\right\rangle\right|\leq M_{f}\|\mathbf{u}\|_{x}^{3}.$
For a univariate, three times differentiable function $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$, $(M_{f},\nu)$-generalized self-concordance [^3] For self-concordance, $\nu=3$. condition is
$|f^{\prime\prime\prime}(x)|\leq M_{f}\left(f^{\prime\prime}(x)\right)^{\frac{\nu}{2}}.$
By sun2019generalized, the composition of a generalized self-concordant function with a linear map is still generalized self-concordant. Hence, it suffices that we show that the functions $f(X)=-\log\det(X)$ and $g(X)=\operatorname{Tr}\left(X^{-p}\right)$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$, $p>0$, are generalized self-concordant for $X$ in the PD cone.
Notice that $f$ is the logarithmic barrier for the PSD cone which is known to be self-concordant (nesterov1994interior). Convergence of Frank-Wolfe for the (D-Opt) Problem is therefore guaranteed. For convergence for the (GTI-Opt) Problem and (A-Opt) Problem, we can show that $g$ is self-concordant on a part of the PD cone.
Theorem 4.
The function $g(X)=\operatorname{Tr}\left(X^{-p}\right)$, with $p>0$, is $\left(3,\frac{(p+2)\sqrt[4]{\alpha^{2p}n}}{\sqrt{p(p+1)}}\right)$-generalized self concordant on $\mathcal{D}=\left\{X\in\mathbb{S}_{++}^{n}\mid 0\prec X\preccurlyeq\alpha I\right\}$.
Proof.
As with the convexity proofs in Section˜A.2, we bring our function back to the univariate case. We define $M(t)=V+tU$ where $V\in D$, $U$ a symmetric matrix and $t\in{\mathbb{R}}$ such that $V+tU\in\mathcal{D}$. By Lemma˜10, the second and third derivatives of $h(t)=\operatorname{Tr}\left(M(t)^{-p}\right)$ are given by
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}\right)$
$h^{\prime\prime\prime}(t)_{\restriction t=0}$ $=-p(p+1)(p+2)\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right).$
Thus, the condition we want to satisfy is
$\left|p(p+1)(p+2)\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right)\right|$ $\leq M_{f}p(p+1)Tr\left(V^{-p}UV^{-1}UV^{-1}\right)^{\nu/2}.$
If the LHS is equal to 0, we are done. Otherwise, we divide both sides by the LHS and first try to lower bound the following fraction.
$\frac{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-\frac{p+1}{2}}\right)^{\nu/2}}{\left|\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-1}UV^{-\frac{p+1}{2}}\right)\right|}$
For that we define $H=V^{-1/2}UV^{-1/2}$ and $L=V^{-p/2}H$.
$\frac{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-\frac{p+1}{2}}\right)^{\nu/2}}{\left|\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-1}UV^{-\frac{p+1}{2}}\right)\right|}$ $=\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\left|\operatorname{Tr}\left(LHL^{\intercal}\right)\right|}$
Using Cauchy-Schwartz, we find
$\left|\operatorname{Tr}\left(LHL^{\intercal}\right)\right|$ $=\left|\left\langle L^{\intercal}L,H\right\rangle\right|$
$\leq\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}.$
$\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\left|\operatorname{Tr}\left(LHL^{\intercal}\right)\right|}$ $\geq\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}}$
$\operatorname{Tr}\left(HH\right)$ $=\operatorname{Tr}\left(V^{p/2}V^{-p/2}HHV^{-p/2}V^{p/2}\right)$
$=\operatorname{Tr}\left(V^{p/2}LL^{\intercal}V^{p/2}\right)$
$\leq\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(V^{2p}\right)}$
$\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}}$ $\geq\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(V^{2p}\right)}}}$
$=\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{\left\|V^{p}\right\|_{F}}}$
$\geq\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{\left\|V^{p}\right\|_{2}}\sqrt[4]{n}}$
By the assumption that V is a feasible point, we have $\left\|V\right\|_{2}=\lambda_{\max}(V)\leq\alpha$.
$\geq\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{\alpha^{p}}\sqrt[4]{n}}$
Then,
$\frac{M_{f}(p(p+1))^{\nu/2}}{p(p+1)(p+2)}\frac{\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}\right)^{\nu-2}}{\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right)}$ $\geq\frac{M_{f}(p(p+1))^{\nu/2}}{p(p+1)(p+2)}\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{\alpha^{p}}\sqrt[4]{n}}.$
We can select $\nu=3$ and re-express the GSC condition as
$1$ $\leq\frac{M_{f}(p(p+1))^{1/2}}{(p+2)}\frac{1}{\sqrt[4]{\alpha^{2p}n}}$
$M_{f}$ $\geq\frac{(p+2)\sqrt[4]{\alpha^{2p}n}}{\sqrt{p(p+1)}}.$
Thus, $g$ is $\left(3,\frac{(p+2)\sqrt[4]{\alpha^{2p}n}}{\sqrt{p(p+1)}}\right)$-generalized self concordant. ∎
Remark.
We conjecture that there is another value of $\nu$ with a tighter value for the constant $M_{f}$ satisfying the condition in Definition˜4.
Corollary 1.
The function $g(X)=\operatorname{Tr}\left(X^{-1}\right)$ is $\left(3,\frac{3\sqrt[4]{\alpha^{2}n}}{\sqrt{2}}\right)$-generalized self concordant on $\mathcal{D}=\left\{X\mid 0\prec X\preccurlyeq\alpha I\right\}$.
Hence, the objectives of the A-Opt and GTI-Opt Problems are generalized self-concordant if we can upperbound the maximum eigenvalue of the information matrix $X(\mathbf{x})$.
Lemma 1.
Let $x\in\mathcal{P}$. Then,
$\lambda_{\max}(X(\mathbf{x}))\leq\max_{i}u_{i}\max_{j}\left\|\mathbf{v}_{j}\right\|_{2}^{2}$
Proof.
$\lambda_{\max}(X(\mathbf{x}))$ $=\lambda_{\max}\left(\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}\right)$
By the Courant Fischer Min-Max Theorem, we have
$\leq\sum\limits_{i=1}^{m}\lambda_{\max}\left(x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}\right).$
Using $\mathbf{x}\leq\mathbf{u}$, yields
$\leq\max_{i}u_{i}\max_{j}\left\|\mathbf{v}_{j}\right\|_{2}^{2}.$
The proof of Theorem˜4 does not translate to the log variant of the A-criterion. Our computational experiments strongly suggest that the function $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ is generalized self-concordant.
Conjecture 1.
The function $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ is $\left(\nu,M\right)$-generalized self concordant for some $\nu\in[2,3]$ and $M>0$ on some subset of PD cone.
Thus, we have convergence of Frank-Wolfe for the D-Opt, A-Opt and GTI-Opt Problems. For the logA-Opt and logGTI-Opt Problems, we have strong empirical evidence that convergence holds but the theoretical proof is still open. Lastly, we aim to show that we have improved convergence rates of Frank-Wolfe for our problems of interest.
3.3 Sharpness
The classical convergence speed of Frank-Wolfe methods given a convex objective is $\Theta(1/\epsilon)$ where $\epsilon$ is the additive primal error to the true optimum $x^{*}$, i.e. $f(\mathbf{x})-f(\mathbf{x}^{*})\leq\epsilon$. For the objectives of the problems defined in Section˜2.1, we can prove a stronger property than convexity, namely sharpness on (a subset of) the feasible region $\mathcal{P}$. This improves the convergence rates of Frank-Wolfe, for the proof see Section˜C and zhao2025new.
Definition 5 (Sharpness (braun2022conditional, Section 3.1.5)).
Let $\mathcal{X}$ be a compact convex set. A convex function $f$ is $(c,\theta)$-sharp (over $\mathcal{X}$) for $0<M<\infty$, $0<\theta<1$ if for all $x\in\mathcal{X}$ and $\mathbf{x}^{*}\in\Omega^{*}_{\mathcal{X}}$, we have
$M\left(f(\mathbf{x})-f(\mathbf{x}^{*})\right)^{\theta}\geq\min_{\mathbf{y}\in\Omega^{*}_{\mathcal{X}}}\left\|\mathbf{x}-\mathbf{y}\right\|$
where $\Omega^{*}_{\mathcal{X}}$ denotes the set of minimizers $\min_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$.
Theorem 5.
The objective functions associated with the Optimal Problem and Fusion Problem
$f(\mathbf{x})$ $=-\log\det(X(\mathbf{x}))$
$f_{F}(\mathbf{x})$ $=-\log\det(X_{C}(\mathbf{x}))$
$g(\mathbf{x})$ $=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$
$g_{F}(\mathbf{x})$ $=\operatorname{Tr}\left(X_{C}(\mathbf{x})^{-p}\right)$
$k(\mathbf{x})$ $=\log\left(\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)\right)$
$k_{F}(\mathbf{x})$ $=\log\left(\operatorname{Tr}\left(X_{C}(\mathbf{x})^{-p}\right)\right)$
with $p>0$ are $(c,1/2)$-sharp for some $c>0$ on $\mathcal{W}:=\operatorname{conv}\left\{x\in\mathcal{P}\cap D\cap{\mathbb{N}}_{0}^{m}\right\}$.
To prove Theorem˜5, we first prove that the functions $f(X)=-\log\det(X)$, $g(X)=\operatorname{Tr}(X^{-p})$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ with $p>0$, are strongly convex on a part of the PD cone.
Definition 6 (Strong Convexity).
A differentiable function $f:D\rightarrow{\mathbb{R}}$ with $D\subseteq{\mathbb{R}}^{n}$ is called $\mu$-strongly convex iff:
$\left\langle\nabla f(\mathbf{x})-\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle\geq\mu\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}\;\;\forall\mathbf{x},\mathbf{y}\in D,$
with $\mu>0$. If $f$ is twice differentiable, an equivalent requirement is:
$\nabla^{2}f(\mathbf{x})\succcurlyeq\mu I\;\;\forall\mathbf{x}\in D.$
That is, $\mu$ is a lower bound on the minimum eigenvalue of the Hessian of $f$.
Lemma 2.
The function $f(X):=-\log\det(X)$ is $\mu$-strongly convex on $\mathcal{D}:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\lambda_{\max}(X)\leq\alpha\}$ with $\mu\leq\frac{1}{a^{2}}$.
Proof.
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma˜9, $f(X)$ is convex and the second derivative of $h$ is given by
$h^{\prime\prime}(t)$ $=\sum\limits_{i=1}^{n}\frac{\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}}{1+t\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)}$
It is sufficient to show that $h^{\prime\prime}(t)\restriction_{t=0}\geq\mu$ to argue strong convexity for $f$.
$h^{\prime\prime}(t)\restriction_{t=0}$ $=\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}$
By using a similar idea to Lemma˜16, we find
$\geq\sum\limits_{i=1}^{n}\lambda_{\min}\left(A^{-1}\right)^{2}\lambda_{i}(B)^{2}$
$=\lambda_{\min}\left(A^{-1}\right)^{2}\operatorname{Tr}\left(B^{2}\right)$
$=\lambda_{\min}\left(A^{-1}\right)^{2}\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{1}{\lambda_{\max}(A)^{2}}$
$\geq\frac{1}{\alpha^{2}}$
$=\mu$
Thus, the function is strongly convex at any point $A\in D$. ∎
Lemma 3.
For $p\in{\mathbb{R}}_{>0}$, the function $g(X):=\operatorname{Tr}\left(X^{-p}\right)$ is $\mu$-strongly convex on the domain $\mathcal{D}:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\lambda_{\max}(X)\leq\alpha\}$ with
$\mu\leq\frac{p(p+1)}{a^{p+2}}.$
Proof.
Let $A\in\mathcal{D}$ and $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. For the computation of $h^{\prime\prime}(t)$, see Lemma˜10.
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}\right).$
Since $A^{-p}$ is positive definite by Corollary˜2 and the trace is cyclic, we have
$=p(p+1)\operatorname{Tr}\left(A^{-\frac{p+2}{2}}BBA^{-\frac{p+2}{2}}\right)$
Note that $BB$ is postive semi-definite. Thus, we can use the bound in fang1994inequalities.
$\geq p(p+1)\lambda_{\min}\left(A^{-(p+2)}\right)\operatorname{Tr}(BB)$
$=p(p+1)\lambda_{\min}\left(A^{-(p+2)}\right)\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{p(p+1)}{\lambda_{\max}\left(A^{p+2}\right)}$
$\geq\frac{p(p+1)}{\alpha^{p+2}}$
Thus, $h(t)$ is strongly convex with $\mu=\frac{p(p+1)}{\alpha^{p+2}}$. ∎
Lemma 4.
For $p\in{\mathbb{R}}_{>0}$, , the function $k(X)=\log\left(\operatorname{Tr}(X^{-p})\right)$ is strongly convex on $\mathcal{D}:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\lambda_{\max}(X)\leq\alpha,\kappa(X)\leq\kappa\}$.
Proof.
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathcal{D}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma˜11, we have
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p\frac{(p+1)\operatorname{Tr}\left(A^{-p-2}BB\right)\operatorname{Tr}\left(A^{-p}\right)-p\operatorname{Tr}\left(A^{-p-1}B\right)^{2}}{\operatorname{Tr}\left(A^{-p}\right)^{2}}.$
Using the Cauchy Schwartz inequality $\operatorname{Tr}\left(XY^{\intercal}\right)^{2}\leq\operatorname{Tr}(XX^{\intercal})\operatorname{Tr}(YY^{\intercal})$ and setting $X=A^{-p/2}$ and $Y=A^{-\frac{p+2}{2}}B$, we have
$\geq p\frac{(p+1-p)\operatorname{Tr}\left(A^{-p}\right)\operatorname{Tr}\left(A^{-p-2}BB\right)}{\operatorname{Tr}(A^{-p})^{2}}$
$=p\frac{\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-p-2}BB\right)}{\operatorname{Tr}\left(A^{-p}\right)}.$
Note that the matrix $BB$ is PSD. The proof of Lemma˜16 can be adapted for one of matrices being symmetric and the other positive definite.
$\geq p\frac{\lambda_{\min}\left(A^{-p-2}\right)\operatorname{Tr}(BB)}{\operatorname{Tr}\left(A^{-p}\right)}$
Observe that $Tr(BB)=\left\|B\right\|_{F}^{2}$. Without loss of generality, we can assume $\left\|B\right\|_{F}=1$.
$\geq p\frac{\lambda_{\min}\left(A^{-p-2}\right)}{n\lambda_{\max}(A^{-p})}$
$=\frac{p}{n\kappa(A)^{p}\lambda_{\max}(A)^{2}}$
$\geq\frac{p}{n\kappa^{p}a^{2}}=\mu$
Thus, $k$ is strongly convex for any $A\in\mathcal{D}$. ∎
By Lemma˜1, we know that the largest eigenvalue of the information matrix can be bounded. As for the minimum eigenvalue, it can be lower bounded on
$\mathcal{W}:=\operatorname{conv}\left\{x\in\mathcal{P}\cap D\cap{\mathbb{N}}_{0}^{m}\right\}$
so the convex hull of all domain-feasible integer points, see Figure˜1. Any point $Y\in\mathcal{W}$ can be represented as a convex combination of the integer vertices $\{V_{i}\}\in\mathcal{W}$. By Lemma˜15, we have
$\lambda_{\min}(Y)$ $=\lambda_{\min}\left(\sum\limits_{i=1}^{n}\gamma_{i}V_{i}\right)$
$\geq\sum\limits_{i=1}^{n}\gamma_{i}\lambda_{\min}(V_{i})$
$\geq\min_{V\in\mathcal{W}}\lambda_{\min}(V)$
$\mathcal{W}$$\mathcal{P}$$D$
Figure 1: A schematic representation of the feasible region $\mathcal{P}$, the domain of the objective $\mathcal{D}$ and the convex hull of vertices that are both feasible and in the domain denoted as $\mathcal{W}$.
Proof of Theorem˜5.
The functions of interest are compositions of strongly convex functions, see Lemma˜2 and Lemma˜3, and linear maps. Observe that the linear maps $X(\mathbf{x})$ and $X_{C}(\mathbf{x})$ are, in general, not injective and hence strong convexity is not preserved.
We can, however, follow the argumentation in braun2022conditional. All of the functions of interest have the form $k(\mathbf{x})=r(B\mathbf{x})$ where $r$ is a strongly convex function. Notice that $B=A^{\intercal}*A^{\intercal}$ where * denotes the column-wise Kroenecker product. Observe that
$\mathcal{W}:=\operatorname{conv}\left(D\cap\mathcal{P}\cap{\mathbb{Z}}^{m}\right)$
is the convex hull of the feasible integer points that are also in the domain. This is a polytope and we may use the bound in braun2022conditional. Notice that the gradient of $k$ is given by $\nabla k(\mathbf{x})=B^{\intercal}\nabla r(B\mathbf{x})$. As in Definition˜5 let $\Omega^{*}_{\mathcal{W}}$ denote the set of minimizers of $\min_{\mathbf{x}\in\mathcal{W}}k(\mathbf{x})$ and let $\mathbf{x}*\in\Omega^{*}_{\mathcal{W}}$. Then, we have $\left\langle\nabla k(\mathbf{x}*),\mathbf{x}-\mathbf{x}*\right\rangle\geq 0$ for $\mathbf{x}\in\mathcal{W}$.
$k(\mathbf{x})-k(\mathbf{x}^{*})$ $=r(B\mathbf{x})-r(B\mathbf{x}^{*})$
Since $r$ is $\mu$-strongly convex we have
$\geq\left\langle\nabla r(B\mathbf{x}^{*}),B\mathbf{x}-B\mathbf{x}^{*}\right\rangle+\mu\frac{\left\|B\mathbf{x}-B\mathbf{x}^{*}\right\|^{2}}{2}$
$=\left\langle\nabla k(\mathbf{x}),\mathbf{x}-\mathbf{x}^{*}\right\rangle+\mu\frac{\left\|B\mathbf{x}-B\mathbf{x}^{*}\right\|_{2}}{2}$
$=\mu\frac{\left\|B\mathbf{x}-B\mathbf{x}^{*}\right\|^{2}}{2}$
Since our region of interest $\mathcal{W}$ is a polytope braun2022conditional holds and there exists a constant $c>0$ such that the distance $\left\|B\mathbf{x}-B\mathbf{y}\right\|$ between any two points $\mathbf{x},\mathbf{y}\in\mathcal{W}$ is bounded by
$\left\|B\mathbf{x}-B\mathbf{y}\right\|$ $\geq c\,\text{dist}\left(\mathbf{x},\{\mathbf{w}\in\mathcal{W}\mid B\mathbf{w}=B\mathbf{y}\}\right).$
In conclusion, we find
$k(\mathbf{x})-k(\mathbf{x}^{*})$ $\geq\mu\frac{\left\|B\mathbf{x}-B\mathbf{x}^{*}\right\|^{2}}{2}\geq\frac{\mu}{2}c^{2}\min_{\mathbf{z}\in\Omega^{*}_{\mathcal{W}}}\left\|\mathbf{x}-\mathbf{z}\right\|^{2}$
Thus, functions are $\left(\frac{\sqrt{2}}{c\sqrt{\mu}},\frac{1}{2}\right)$-sharp. ∎
Regarding the constant $c$ in the previous proof, we find
$\|B\mathbf{x}-B\mathbf{y}\|_{F}$ $=\|A^{\intercal}\operatorname{diag}(\mathbf{x})A-A^{\intercal}\operatorname{diag}(y)A\|_{F}$
$=\|A^{\intercal}\operatorname{diag}(\mathbf{x}-\mathbf{y})A\|_{F}.$
Let $C=A^{\intercal}$ and $B=\operatorname{diag}(\mathbf{x}-\mathbf{y})A$ and let us use the bound in Lemma˜17.
$\geq\sigma_{\min}(A)\|\operatorname{diag}(\mathbf{x}-\mathbf{y})A\|_{F}$
$\geq\sigma_{\min}(A)^{2}\|\operatorname{diag}(\mathbf{x}-\mathbf{y})\|_{F}$
$=\sigma_{\min}(A)^{2}\|\mathbf{x}-\mathbf{y}\|_{2}$
Note that $\|\mathbf{x}-\mathbf{y}\|_{2}\geq\text{dist}\left(\mathbf{x},\{\mathbf{w}\in\mathcal{W}\mid B\mathbf{w}=B\mathbf{y}\}\right)$ and thus, we have found $c=\sigma_{min}(A)^{2}$. Note that since $A$ has full column rank, the minimum singular value is strictly greater than 0.
We have thus established sharpness for all objective functions on $\mathcal{W}$. Observe that $\mathcal{W}=\mathcal{P}$ for the Fusion Problems, thus, linear convergence of Frank-Wolfe is guaranteed on them. In general, however, we have $\mathcal{P}\backslash\mathcal{W}\cap D\neq\emptyset$, see also Figure˜1. That is there exists domain feasible points $\mathbf{y}\in\mathcal{P}$ which cannot be expressed as the convex combination of domain feasible integer points in $\mathcal{P}$. Hence, we have only linear convergence of Frank-Wolfe on the continuous subproblems if the corresponding continuous solution is in $\mathcal{W}$. In this case, we may use Theorem˜6 in Section˜C to argue linear convergence. For the Equation˜D-Opt Problem, we could adapt the proof in zhao2025new to the Blended Pairwise Conditional Gradient (BPCG) to argue linear convergence on all the subproblems, see Proposition˜1 in Section˜C. For the A-criterion and its log variant, we have only linear convergence on nodes where the continuous solution is in $\mathcal{W}$. We cannot guarantee that this is the case for all subproblems at the nodes.
(a) AA-criterion
(b) $D$-criterion
(c) $\log A$-criterion
Figure 2: Trajectory of the dual gap at the root node compared to $1/2t$, $1/t^{2}$ and $1/e^{ct}$.
From our experiments, we have empirical evidence that we have linear convergence of the dual gap for all Optimal Problems. In Figure˜2, we show the trajectory of the dual gap at the root node for the $A$-criterion, $D$-criterion and $\log A$-criterion. After a burn-in phase, we see that the trajectory becomes parallel to $1/e^{ct}$ for some constant $c>0$ which is a representation of the linear convergence.
In summary, we have definitely accelerated convergence rates for the Fusion Problems. In case of the Optimal Problems, it will depend on the position of the continuous solution, so the input experiment data. In our own experiments, we could observe accelerated rates.
Objective
Convergence
Linear
Convergence
$f_{F}(\mathbf{x})=\log\det(X_{C}(\mathbf{x}))$ $\checkmark$ $\checkmark$
$g_{F}(\mathbf{x})=\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})$ $\checkmark$ $\checkmark$
$k_{F}(\mathbf{x})=\log\left(\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})\right)$ $\checkmark$ $\checkmark$
$f(\mathbf{x})=-\log\det(X(\mathbf{x}))$ $\checkmark$ $\checkmark$
$g(\mathbf{x})=\operatorname{Tr}(X(\mathbf{x})^{-p})$ $\checkmark$ Not guaranteed
$k(\mathbf{x})=\log\left(\operatorname{Tr}(X(\mathbf{x})^{-p})\right)$ Conjecture Not guaranteed
Table 1: Convergence of Frank-Wolfe on different objective functions.
4 Solution methods
The main goal of this paper is to propose a new solution method for the Optimal and Fusion Problems under the A-criterion and D-criterion based on the novel framework Boscia.jl and assess its performance compared to several other convex MINLP approaches. In the following, we introduce the chosen MINLP solvers and state the necessary conditions and possible reformulations that are needed.
Branch-and-bound with Frank-Wolfe algorithms (Boscia).
The new framework introduced in hendrych2023convex is implemented in the Julia package Boscia.jl. It is a Branch-and-Bound (BnB) framework that utilizes Frank-Wolfe methods to solve the relaxations at the node level. The Frank-Wolfe algorithm (frank1956algorithm; braun2022conditional), also called Conditional Gradient algorithm (levitin1966constrained), and its variants are first-order methods solving problems of the type:
$\min_{\mathbf{x}\in\mathcal{X}}\;$ $f(\mathbf{x})$
where $f$ is a convex, Lipschitz-smooth function and $\mathcal{X}$ is a compact convex set. These methods are especially useful if linear minimization problems over $\mathcal{X}$ can be solved efficiently. The Frank-Wolfe methods used in Boscia.jl are implemented in the Julia package FrankWolfe.jl, see besanccon2022frankwolfe; besanccon2025improved.
At each iteration $t$, Frank-Wolfe solves the linear minimization problem over $\mathcal{X}$ taking the current gradient as the linear objective, resulting in a vertex $v_{t}$ of $\mathcal{X}$. The next iterate $x_{t+1}$ is computed as a convex combination of the current iterate $x_{t}$ and the vertex $v_{t}$. Many Frank-Wolfe variants explicitly store the vertex decomposition of the iterate, henceforth called the active set. We utilize the active set representation to facilitate warm starts in Boscia by splitting the active set when branching.
One novel aspect of Boscia.jl is its use of a Bounded Mixed-Integer Linear Minimization Oracle (BLMO) as the Linear Minimization Oracle (LMO) in Frank-Wolfe. Typically, the BLMO is a MIP solver but it can also be a combinatorial solver. This leads to more expensive node evaluations but has the benefit that feasible integer points are found from the root node and the feasible region is much tighter than the continuous relaxation for many problems. In addition, Frank-Wolfe methods can be lazified, i.e. calling the LMO at each iteration in the node evaluation can be avoided, see (braun2017lazifying).
In the case of OEDP, strong lazification is not necessary since the corresponding BLMO is very simple. The feasible region $\mathcal{P}$ is just the scaled probability simplex intersected with integer bounds, which is amenable to efficient linear optimization. Given a linear objective $d$, we first assign $\mathbf{x}=\mathbf{l}$ to ensure that the lower bound constraints are met. Next, we traverse the objective entries in increasing order, adding to the corresponding variable the value of $\max\{u_{i}-l_{i},N-\sum(\mathbf{x})\}$. This way, we ensure that both the upper-bound constraints and the knapsack constraint are satisfied. The LMO over the feasible set can also be cast as a simple network flow problem with $m$ input nodes connected to a single output node, which must receive a flow of $N$ while the edges respect the lower and upper bounds. The scaled probability simplex also allows for a simple heuristic taking a fractional point and outputting an integer feasible point on the probability simplex.
Due to the convexity of the objective, the difference $\left\langle\nabla f,x_{t}-v_{t}\right\rangle$ is upper bounding the primal gap $f(x_{t})-f(x^{*})$ at each iteration. We call the quantity the dual gap (or the Frank-Wolfe gap). The dual gap can therefore be used as a stopping criterion. Frank-Wolfe’s error adaptiveness can be exploited to a) solve nodes with smaller depth with a coarser precision and b) dynamically stop a node evaluation if the lower bound on the node solution exceeds the incumbent.
Observe that in contrast to the epigraph-based formulation approaches that generate many hyperplanes, this method works with the equivalent of a single supporting hyperplane given by the current gradient and moves this hyperplane until it achieves optimality. It is known that once the optimal solution is found, a single supporting hyperplane can be sufficient to prove optimality (described e.g. for generalized Benders in sahinidis1991convergence). Finding this final hyperplane, however, may require adding many hyperplanes beforehand at suboptimal iterations. In the case of the problems discussed in this paper, the constraint polytope is uni-modular. Adding hyperplanes created from the gradient will not maintain this structure and in consequence, yields a numerically more challenging MIP. Our approach, on the other hand, keeps the polytope and thereby its uni-modularity intact.
A new development during this work was the implementation of a domain oracle in Boscia.jl. As stated earlier, not all points in the feasible region are domain feasible for the objective functions. Thus, we cannot start Boscia (and Frank-Wolfe) at a random point. We therefore added the option of a domain oracle and a domain point function. The former checks whether a given point is domain feasible, in our case by checking if the associated information matrix is regular. The latter provides a domain feasible point given the feasible region and additional bound constraints from the node. The returned point has to respect the global constraints as well as the node specific constraints. The domain point function is called during branching if the splitting of the active set results in an invalid point. A small projection problem is then solved, also using Frank-Wolfe, to find a new point within the intersection of the feasible region and the domain. Note that we stop the Frank-Wolfe method once we are in the intersection.
For the line search problem within Frank-Wolfe, we utilize the Secant line search which has shown a superior performance for generalized self concordant functions, see hendrych2025secant.
Outer approximation (SCIP OA).
Outer Approximation schemes are a popular and well-established way of solving MINLPs (kronqvist2019review). This approach requires an epigraph formulation of (OEDP):
$\begin{split}\min_{t,\mathbf{x}}\;&t\\ \text{s.t. }&t\geq\log\left(\phi_{p}(\mathbf{x})\right)\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$ (E-OEDP)
This approach approximates the feasible region of (E-OEDP) with linear cuts derived from the gradient of the non-linear constraints, in our case $\nabla f$. Note that this requires the information matrix $X(\mathbf{x})$ at the current iterate $\mathbf{x}$ to be positive definite, otherwise an evaluation of the gradient is not possible or rather it will evaluate to $\infty$. The implementation is done with the Julia wrapper of SCIP, (bestuzheva2021scip; bestuzheva2023enabling). Note that generating cuts that prohibit points leading to singular $X(\mathbf{x})$, we will refer to them as domain cuts, is non-trivial. Thus, we are applying this approach to the Fusion Problems only since their information matrix is always positive definite.
LP/NLP branch-and-bound with direct conic formulation (Direct Conic).
Another Outer Approximation approach, as implemented in Pajarito.jl (CoeyLubinVielma2020), represents the non-linearities as conic constraints. This is particularly convenient in combination with the conic interior point solver Hypatia.jl (coey2022solving) as it implements the $\log\det$ cone (the epigraph of the perspective function of $\log\det$) directly.
$\mathcal{K}_{\log\det}:=\text{cl}\left\{(u,v,W)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\mathbb{S}_{++}^{n}\mid u\leq v\log\det(W/v)\right\}$
The formulation of (D-Opt) then becomes
$\begin{split}\max_{t,\mathbf{x}}\;&t\\ \text{s.t. }&(t,1,X(\mathbf{x}))\in\mathcal{K}_{\log\det}\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$
For the representation of the trace inverse, we utilize the dual of the separable spectral function cone (coey2022performance, Section 6):
$\mathcal{K}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathcal{Q})\mid u\geq v\varphi(w/v)\right\}$
For our purposes, $\mathcal{Q}$ is the PSD cone and the spectral function $\varphi$ is the negative square root whose convex conjugate is precisely the trace inverse, see (coey2023conic, Table 1) and (pilanci2015sparse, Collary 1) [^4] For further details, see this discussion formulationforum..
$\mathcal{K}^{*}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathbb{S}_{+}^{n})\mid v\geq u/4\operatorname{Tr}((w/u)^{-1})\right\}$
The conic formulation of the (A-Opt) is therefore
$\begin{split}\min_{t,\mathbf{x}}\;&4t\\ \text{s.t. }&(1,t,X(x))\in\mathcal{K}^{*}_{\text{sepspec}}\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$
Additionally, the conic formulation allows the computation of domain cuts for $\mathbf{x}$. Hence, this solver can be used on all problems. Note that we use HiGHS (huangfu2018parallelizing) as a MIP solver within Pajarito.jl.
LP/NLP branch-and-bound with second-order cone formulation (SOCP).
The conic solver from the previous approach was developed fairly recently. A previous conic approach was introduced first in sagnol2011computing for the continuous case and later extended to the mixed-integer case in sagnol2015computing. They showed OEDP under the A-criterion and D-criterion can be formulated as a second-order cone program (SOCP). The SOCP model for (D-Opt) is given by
$\begin{split}\max_{\mathbf{w},Z,T,J}\;&\prod\limits_{j=1}^{n}\left(J_{jj}\right)^{\frac{1}{n}}\\ \text{s.t. }&\sum\limits_{i=1}^{m}A_{i}Z_{i}=J\\ &J\text{ is a lower triangle matrix}\\ &\left\|Z_{i}\mathbf{e}_{j}\right\|^{2}\leq T_{ij}\mathbf{w}_{i}\quad\forall i\in[m],j\in[n]\\ &\sum\limits_{i=1}^{m}T_{ij}\leq J{jj}\quad\forall j\in[n]\\ &T_{ij}\geq 0\quad\forall i\in[m],j\in[n]\\ &\sum\limits_{i=1}^{m}\mathbf{w}_{i}=N,\;\mathbf{w}\geq 0\end{split}$
where $J$ is a $n\times n$ matrix, $Z$ and $T$ are $m\times n$ matrices. Observe that the problem size has increased significantly with both the number of variables and number of contraints being $\geq m(2n+1)$.
The SOCP formulation for (A-Opt) is
$\begin{split}\min_{\mathbf{w},\bm{\mu},Y}\;&\sum\limits_{i=1}^{m}\mu_{i}\\ \text{s.t. }&\sum\limits_{i=1}^{m}A_{i}Y_{i}=\sum\limits_{i=1}^{m}\bm{\mu}_{i}\\ &\left\|Y_{j}\right\|^{2}\leq\bm{\mu}_{j}\mathbf{w}_{i}\quad\forall i\in[m]\\ &\bm{\mu}\geq 0,\;\mathbf{w}\geq 0\\ &\sum\limits_{i=1}^{m}\mathbf{w}_{i}=N\end{split}$
where $Y$ is a $m\times n$ matrix and $\bm{\mu}$ is a $m\times 1$ vector. The problem size does not increase as much, with $m(2+n)$ variables and $3m+2$ constraints.
To solve the SOCP formulation, we again utilize Pajarito.jl with HiGHS as a MIP solver and Hypatia.jl as a conic solver. For the terms involving the norm, we make use of Hypatia.jl’s EpiNormEuclCone and EpiPerSquareCone cones which do not require us to deconstruct the norm terms into second order cones.
A custom branch-and-bound for (OEDP) (Co-BnB).
The most general solver strategy for (OEDP) with matrix means criteria was introduced in ahipacsaouglu2021branch. Like Boscia.jl, it is a Branch-and-Bound-based approach with a first-order method to solve the node problems. The first-order method in question is a coordinate-descent-like algorithm. As the termination criterion, this method exploits that the objective function is a matrix mean and shows the connection of the resulting optimization problem
$\begin{split}\max_{\mathbf{w}}\,&\,\log\left(\phi_{p}(X(\mathbf{w}))\right)\\ \text{s.t. }\,&\sum\limits_{i=1}^{m}w_{i}=1\\ &\mathbf{w}\geq 0\\ &w_{i}N\in{\mathbb{Z}}\,\forall i\in[m]\end{split}$ (M-OEDP)
to the generalization of the Minimum Volume Enclosing Ellipsoid Problem (MVEP) (ahipacsaouglu2021branch). The variables $w$ can be interpreted as a probability distribution and the number of times the experiments are to be run is $wN$. Concerning (OEDP), one can say $x=wN$.
Note first that we have reimplemented the method in Julia and modified it so that we solve the formulation as depicted in (OEDP). Secondly, we have improved and adapted the step size rules within the first-order method, see Section˜B. Note further that the solver was developed for instances with a plethora of experiments and very few parameters. The solver employs the simplest Branch-and-Bound tree, i.e. with the most fractional branching rule and traverses the tree with respect to the minimum lower bound. In the next section, we will see that the method works well in cases where $n$ is small but struggles if the value of $n$ increases.
5 Computational experiments
In this section, we present the computational experiments for the Optimal Problem and Fusion Problem, both under the A- and D-criterion, respectively. The resulting problems will be referred to as the A-Fusion Problem (AF), D-Fusion Problem (DF), A-Optimal Problem (AO), and D-Optimal Problem (DO).
Furthermore, we run the Optimal and Fusion Problems under the GTI-criterion for different values of $p$ using permissible solvers. Lastly, we showcase the performance of Boscia.jl and Co-BnB on problems of large dimension, i.e. number of variables between 300 and 500.
Experimental setup.
For the instance generation, we choose the number of experiments $m\in\{50,60,80,\linebreak 100,120\}$, the number of parameters $n\in\{\lfloor m/4\rfloor,\lfloor m/10\rfloor\}$, and the number of allowed experiments $N=\lfloor 1.5n\rfloor$ for the Optimal Problems and $N\in[m/20,m/3]$ for the Fusion Problems. The lower bounds are zero. Note that for the Fusion Problems, the fixed experiments are encoded in a separate matrix. The upper bounds are randomly sampled between 1 and $N/3$ for (A-Opt) and (D-Opt). In the Fusion case, they are sampled between 1 and $m/10$. We generate both independent and correlated experiment data. Also, note that the matrices generated are dense. Seeds used are $[1,2,3,4,5]$. In total, there are 50 instances for each combination of problem and data set.
Experiments were run on a cluster equipped with Intel Xeon Gold 6338 CPUs running at 2 GHz and a one-hour time limit. The Julia version used is 1.10.2, Boscia.jl v0.1.35, FrankWolfe.jl v0.4.13, HiGHS.jl v1.15.0, Hypatia.jl v0.7.4, Pajarito.jl v0.8.2, SCIP.jl v0.12.3. Note that we use the unregistered package PajaritoExtras.jl to model the dual of the separable spectral function cone. The source code is hosted on GitHub [^5] https://github.com/ZIB-IOL/OptimalDesignWithBoscia.
Start solution.
Note that both the objectives (D-Opt) and (A-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank. This is the case for the Fusion Problem, not necessarily for the Optimal Problem. Both Boscia.jl and Co-BnB require a feasible starting point $\mathbf{z}_{0}$. For its construction, we find a set $S\subset[m]$ of $n$ linearly independent experiments, i.e. $n$ linearly independent rows of $A$. Assign those experiments their upper bound. If the sum $\sum\mathbf{z}_{0}$ exceeds $N$, remove 1 from the experiment with the largest entry. If the sum is less than $N$, pick an experiment in $\S[m]\backslash$ S at random and assign it as many runs as possible. Repeat until the sum is equal to $N$. Note that due to the monotonic progress of both first-order methods, the current iterate will never become domain infeasible, i.e. singular.
Results.
An overview of the results of the computational experiments is given in Table˜2 and extensive results can be found in Section˜D. The new framework Boscia.jl shows a superior performance compared to the other methods both in terms of time and number of instances solved to optimality. In comparison to the Outer Approximation methods, for example, it solves nearly twice as many instances.
Co-BnB is the only competitive method. It mostly solves fewer instances than Boscia.jl, the exception being the A-Fusion Problem with both data sets. With independent data, Co-BnB is faster solving the same number of instances. In case of the correlated data set, Co-BnB solves roughly twice as many instances as Boscia.jl. In Table˜5, we see that Boscia.jl and Co-BnB have similar magnitudes in number of nodes for both A-Fusion Problems whereas for all other problems Boscia.jl has significantly fewer nodes than Co-BnB. As stated previously, node evaluation in Boscia.jl is more computationally expensive. The fact that the magnitudes of nodes is similar points to the optimal solution being strictly in the interior. Note that in general, Co-BnB fares well for the instances where $n=\lfloor m/10\rfloor$ as it was designed with such problems. It struggles for the instances where $n=\lfloor m/4\rfloor$. In Figure˜3, we see that the absolute gap of Co-BnB increasing drastically as the fraction between number of parameters and number experiments grow larger. While Boscia.jl also shows an increase, it is much more moderate. Co-BnB implements no advanced Branch-and-Bound tree specifications like a better traverse strategy or branching strategy. A greater value of $N$ naturally increases the size of the tree and the number of nodes to be processed. Boscia.jl has the advantage here since it finds many integer feasible points while solving the relaxations which have the potential to improve the incumbent. A better incumbent, in turn, lets us prune non-improving nodes early on. Overall, fewer nodes have to be investigated by Boscia.jl.
(a) A-Optimal Problem with independent data
(b) D-Optimal Problem with independent data
(c) A-Optimal Problem with correlated data
(d) D-Optimal Problem with correlated data
Figure 3: Geometric mean and geometric standard deviation of the absolute gap between the incumbent and lower bound over the conditioning of the problem.
In terms of time, Boscia.jl also shows promising results, especially for instances of larger scale. For small-scale instances, the Outer Approximation approaches are fast, in some cases faster than Boscia.jl. This can be observed in Figures˜4 and 5. The plots also highlight that the Outer Approximation Schemes solve fewer instances than the two Branch-and-Bound Approaches. It should be noted that SCIP + OA simply timed out often. On the other hand, the MIP solvers for the conic formulations with Pajarito.jl often struggled to solve the resulting MIP problems. SOCP, for example, only solved about a quarter of the instances of the D-Fusion Problem and D-Optimal Problem. In addition, the many solutions provided by SOCP in case of a timeout were not always domain feasible, and thus, the original objective function could not be evaluated at the provided solutions. The superiority of the Branch-and-Bound approaches stems from the fact that the relaxations of both of them keep the simple feasible region intact. Whereas the Outer Approximation schemes add many additional constraints, i.e. cuts, resulting in larger LPs to be solved, see Table˜5 for the average number of cuts. Furthermore, these cuts are dense, i.e. have many nonzero entries, leading to an increased difficulty for the LP solvers.
(a) A-Optimal Problem with independent data
(b) D-Optimal Problem with independent data
(c) A-Optimal Problem with correlated data
(d) D-Optimal Problem with correlated data
Figure 4: The number of instances solved to optimality over time for the A-Optimal Problem and D-Optimal Problem with both data sets. The upper plots are with the inpdendent data set, the lower ones with correlated data. Note that SCIP+OA is not applicable on the Optimal Problems.
(a) A-Fusion Problem with independent data
(b) D-Fusion Problem with independent data
(c) A-Fusion Problem with correlated data
(d) D-Fusion Problem with correlated data
Figure 5: The number of instances solved to optimality over time for the A-Fusion Problem and D-Fusion Problem with both data sets. The upper plots are with the independent data set, the lower ones with correlated data.
For the (GTI-Opt) Problem, the results are given in Table˜3. Note that we used Boscia.jl on both the (GTI-Opt) and the (logGTI-Opt), the latter titled as "Boscia Log" in the table. Boscia.jl also shows a better performance here than Co-BnB and SCIP + OA. Interestingly, the (logGTI-Opt) objective seems to lead to a better conditioning of the problem. This further motivates an investigation into generalized self-concordance of those objectives.
In Table˜4, we see the relative and absolute gap of Boscia.jl and Co-BnB for instances with sizes from 300 and 500 variables. Boscia.jl showcases better gap values especially for the problems under the D-criterion. Note that almost none of the instances were solved to optimality.
We performed a hyperparameter search for Boscia.jl the results of which can be found in Tables˜8, 9 and 10.
Aside from the performance comparison of the solvers, we investigate how the problems themselves compare to each other and if the difficulty of the instances is solver-dependent.
It can be observed in Table˜2 that most solvers solve fewer instances under the A-criterion. The notable exception is SOCP on the A-Fusion Problem where it solves more instances to optimality compared to the D-criterion.
Taking a look at some example contour plots shown in Figure˜6, we observe that the contour lines for the A-criterion are steeper than those of the D-criterion for both the Optimal Problem and the Fusion Problem, respectively. This points to the conditioning increasing.
(a) D-Optimal Correlated
(b) A-Optimal Correlated
(c) D-Fusion Independent
(d) A-Fusion Independent
Figure 6: Example contour plots in two dimensions for objectives of both Optimal Problems with correlated data and both Fusion Problems with independent.
In terms of the data, one could assume that all problems would be easier to solve with independent data. Noticeably, this is not the case, rather, it depends on the problem type. The Fusion Problems are easier to solve with independent data, the Optimal Problems are more often solved with the correlated data. Figures 7(a) to 7(d) depict the progress of the incumbent, lower bound, and dual gap within Boscia.jl for selected instances of each combination of problem and data type.
Interestingly, the independent data leads to proof of optimality, i.e. the optimal solution is found early on and the lower bound has to close the gap, regardless of the problem, see Figures˜7(a) and 7(c). The difference in problems has, however, an impact on how fast the lower bound can catch up with the incumbent. In the case of the Optimal Problem there is likely a larger region around the optimal solution where the corresponding points/designs $\mathbf{x}$ provide roughly the same information. These other candidates have to be checked to ensure the optimality of the incumbent and thus the solving process slows down. Note that Boscia.jl can utilize strong convexity and sharpness of the objective to improve the lower bound. However, the improvement via sharpness is currently only applicable if the optimum is unique which is not the case.
On the other hand, the correlated data leads to solution processes that are very incumbent-driven, i.e. most improvement on the dual gap stems from the improving incumbent, not from the lower bound, as seen in Figures˜7(b) and 7(d). Incumbent-driven solution processes can be identified by the dual gap making sudden jumps and the absence of (a lot of) progress between these jumps. As before, the solution process speed depends on the Problem. In Figure˜7(b), the dual gap makes big jumps throughout most of the solving process, in contrast to the dual gap Figure˜7(d). This indicates that the optimal solution of the Fusion Problem is strictly in the interior and thus, will not be found early as a vertex of a relaxation. The key ingredient for improvement in this case will be the incorporation of more sophisticated primal heuristics in Boscia.jl.
Overall, Boscia.jl shows a very promising performance and further research and development will be performed to improve the performance even further.
(a) D-Optimal Problem with independent data and $n=10$
(b) D-Optimal Problem with correlated data and $n=10$
(c) A-Fusion Problem with independent data and $n=10$
(d) A-Fusion Problem with correlated data and $n=25$
Figure 7: Progress of the incumbent and the lower bound on the left and progress of the absolute gap on the right. All instances have 100 variables.
Table 2: Comparing the performance of Boscia.jl, Pajarito.jl, SCIP, Co-BnB and SOCP on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data. Shown are the percentage of solved instances and the geometric mean of the solve time (shifted by 1 second) over all instances. Note that there are 50 instances for each problem and data set.
For more extensive results like number of cuts/nodes, see Table˜5 in Section˜D.
Boscia Co-BnB Direct Conic SOCP SCIP OA
Type Corr. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
AO no 52 % 374.74 48 % 401.67 8 % 2382.52 20 % 1578.03 - -
AO yes 68 % 227.69 60 % 301.8 20 % 1515.23 20 % 1861.35 - -
AF no 80 % 62.83 80 % 50.92 4 % 2846.5 26 % 1686.2 38 % 463.65
AF yes 26 % 1368.83 52 % 311.36 24 % 1148.87 20 % 2131.81 14 % 1480.73
DO no 68 % 136.92 58 % 216.31 20 % 1241.37 20 % 2169.4 - -
DO yes 100 % 2.56 70 % 101.89 10 % 1398.01 10 % 2526.74 - -
DF no 94 % 3.49 86 % 22.79 12 % 1906.61 18 % 2213.03 50 % 324.11
DF yes 60 % 54.01 56 % 153.26 6 % 2284.06 6 % 2934.61 28 % 752.15
Table 3: Comparing the performance of Boscia.jl,SCIPand Co-BnB on the General-Trace-Inverse (GTI) Optimal Problem and GTI-Fusion (GTIF) Problem for different values of $p$ and under independent and correlated data. Note the the objective function of Boscia has the form $\phi_{p}(X(\mathbf{x}))=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$. The column titled "Boscia Log" shows the runs using $\log\phi_{p}$ as the objective. Shown are the percentage of solved instances and the geometric mean of the solve time (shifted by 1 second) over all instances. Note that there are 30 instances for each problem and data set.
For extensive results, see Tables˜6 and 7 in Section˜D.
Boscia Boscia Log Co-BnB SCIP+OA
p Corr. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
Optimal Problem
0.25 no 67 % 162.41 87 % 27.13 50 % 359.92 - -
0.25 yes 67 % 110.25 77 % 85.83 60 % 264.22 - -
0.5 no 53 % 246.61 70 % 85.87 50 % 430.08 - -
0.5 yes 60 % 129.86 57 % 245.36 57 % 313.7 - -
0.75 no 37 % 799.73 43 % 505.86 50 % 471.95 - -
0.75 yes 47 % 541.67 43 % 557.15 50 % 357.77 - -
1.5 no 43 % 659.25 40 % 576.51 50 % 619.95 - -
1.5 yes 47 % 418.43 53 % 306.7 47 % 506.3 - -
2.0 no 43 % 661.29 50 % 562.4 50 % 630.18 - -
2.0 yes 43 % 644.53 53 % 297.98 47 % 611.8 - -
Fusion Problem
0.25 no 90 % 6.57 100 % 1.71 50 % 725.52 47 % 376.08
0.25 yes 53 % 138.78 57 % 184.83 27 % 1449.73 27 % 927.92
0.5 no 87 % 9.1 90 % 3.63 47 % 787.45 43 % 349.14
0.5 yes 57 % 175.85 53 % 271.12 30 % 1547.49 20 % 850.41
0.75 no 70 % 85.39 87 % 42.47 57 % 353.3 43 % 415.34
0.75 yes 37 % 741.57 43 % 434.78 27 % 1639.63 23 % 1053.8
1.5 no 77 % 91.83 67 % 129.57 47 % 780.56 40 % 406.94
1.5 yes 37 % 715.74 50 % 214.97 30 % 1415.97 23 % 885.83
2.0 no 73 % 96.28 70 % 117.1 57 % 797.56 40 % 402.28
2.0 yes 33 % 905.59 53 % 202.0 33 % 1305.28 20 % 923.45
Table 4: Comparing Boscia and Co-BnB on large instances. The dimensions are $m=300,400,500$. We showcase only the absolute and relative gap. Note that the majority of the instances were solved to optimality. Boscia solved 3 instances of the D-Fusion Problem with independent data. The relative gap and the absolute gap are computed for the instances on which at least one method did not terminate within the time limit. That means it excludes the instances on which all methods terminated. The "log" prefix indicates that Boscia was used with the $\log\operatorname{Tr}$ objective.
Boscia Co-BnB
Problem Corr. # inst. relative gap absolute gap relative gap absolute gap
A no 12 188.2483 3.8117 2.4036 5.2537
A yes 12 Inf 0.0024 Inf 5.8384
AF no 12 0.2243 0.0019 0.9264 1.0258
AF yes 12 Inf 0.0049 Inf 5.6742
log A no 12 2.4047 5.2608 2.4036 5.2537
log A yes 12 Inf 5.6258 Inf 5.8384
log AF no 12 1.0491 2.6105 0.9264 1.0258
log AF yes 12 0.0955 6.9323 Inf 5.6742
DF no 12 0.0169 0.01 Inf 38.0805
DF yes 12 0.9819 0.4054 Inf 343.1067
D no 12 0.2295 0.1577 Inf 69.0189
D yes 12 0.4642 0.091 673.4528 73.5618
6 Conclusion
We proposed a new approach for the Optimal Experiment Design Problem based on the Boscia.jl framework and proved convergence of the method on the problems. Our approach exhibits a superior performance compared to other MINLP approaches. In addition, it also outperforms the approach specifically developed for the OEDP, in particular for large-scale instances and a larger number of parameters. This superiority can be explained by the fact that Boscia.jl keeps the structure of the problem intact and that it utilizes a combinatorial solver to find integer feasible points at each node.
Acknowledgments
Research reported in this paper was partially supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM,05M20ZBM) and the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+.
A Additional proofs and calculations
Note that the following lemmas and the proofs are not new but they are stated here for completeness.
A.1 General power of a matrix
To start, we present some smaller statements about the power of a matrix. These are used in the consequent proofs.
Lemma 5.
Let $X\in{\mathbb{R}}^{n\times n}$ be an invertible, symmetric matrix. Then, the logarithm $\log(X)$ is symmetric.
Proof.
Let $V$ denote the matrix consisting of eigenvectors of $X$. Observe that $V$ is an orthogonal matrix and w.l.o.g. we assume $V$ is scaled such that it is orthonormal. Next, let $D$ denote the diagonal matrix whose entries are the eigenvalues of $X$. The logarithm of $X$ is then defined as
$\log(X)$ $=V\log(D)V^{-1}$
where $\log(D)$ is the diagonal matrix with logarithms of the eigenvalues. In our case, $V$ is orthonomal, so $V^{-1}=V^{\intercal}$. Hence,
$\log(X)^{\intercal}$ $=\left(V\log(D)V^{\intercal}\right)^{\intercal}=V\log(D)V^{\intercal}=\log(X)$
Lemma 6.
Let $X\in{\mathbb{R}}^{n\times n}$ by an invertible matrix and $r\in{\mathbb{R}}$. Then, the eigenvalues of the power matrix are
$\lambda_{i}(X^{r})=\lambda_{i}(X)^{r}\quad\forall i=[n]$
and the eigenvectors of the power matrix are the same as those of $X$.
Proof.
Let $\mathbf{v}$ be an eigenvector of $X$ and let $\lambda_{\mathbf{v}}$ denote the corresponding eigenvalue.
$X^{r}\mathbf{v}$ $=\left(\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}\right)\mathbf{v}=\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}\mathbf{v}$
For any square matrix $A$ and a natural number $m\in{\mathbb{N}}$, we have that $A^{m}\mathbf{v}=\lambda_{\mathbf{v}}(A)^{m}\mathbf{v}$. Also, the logarithm of a matrix $A$ has the same eigenvectors as the original and its eigenvalues are $\log(\lambda_{i}(A))$.
$=\left(\sum\limits_{k=0}^{\infty}\frac{r^{k}\log(\lambda_{\mathbf{v}}(X))^{k}}{k!}\right)\mathbf{v}=\exp(r\log(\lambda_{\mathbf{v}}(X)))\mathbf{v}$
$=\lambda_{\mathbf{v}}(X)^{r}\mathbf{v}$
This concludes the proof. ∎
Combining Lemma˜5 and Lemma˜6, we get the following corollary.
Corollary 2.
For $r\in{\mathbb{R}}$, if $X\in\mathbb{S}_{++}^{n}$, then $X^{r}\in\mathbb{S}_{++}^{n}$.
Lemma 7.
Let $r\in{\mathbb{R}}$ and let the matrix $X$ be diagonalizable. Let $f(X)=\operatorname{Tr}(X^{r})$. The gradient of $f$ is then
$\nabla f(x)=rX^{r-1}$
Proof.
We use the definition of $X^{r}$ to prove the result.
$\operatorname{Tr}(X^{r})$ $=\operatorname{Tr}\left(\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}\right)=\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\operatorname{Tr}\left(\log(X)^{k}\right)$
Thus, using the sum rule, $(\operatorname{Tr}(X^{n}))^{\prime}=nX^{n-1}$ for any positive integer $n$ and $\nabla\operatorname{Tr}(\log(X))=X^{-1}$, we have
$\nabla\operatorname{Tr}(X^{r})$ $=\sum\limits_{k=1}^{\infty}\frac{r^{k}}{k!}k{\log(X)}^{k-1}X^{-1}$
$=rX^{-1}\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}$
$=rX^{-1}X^{r}=rX^{r-1}$
This concludes the proof. ∎
Lemma 8.
For $X\in\mathbb{S}_{++}^{n}$, we have $X^{-p}=\left(X^{-1}\right)^{p}$.
Proof.
$\left(X^{-1}\right)^{p}$ $=\sum\limits_{k=0}^{\infty}\frac{p^{k}}{k!}\log\left(X^{-1}\right)^{k}$
$\log\left(X^{-1}\right)$ $=V\log(D^{-1})V^{\intercal}=-V\log(D)V^{\intercal}=-\log(X)$
$\left(X^{-1}\right)^{p}$ $=\sum\limits_{k=0}^{\infty}\frac{p^{k}}{k!}\left(-\log(X)\right)^{k}$
$=\sum\limits_{k=0}^{\infty}\frac{(-p)^{k}}{k!}\log(X)^{k}$
$=X^{-p}$
Remark.
On the contrary, $\left(X^{p}\right)^{-1}\neq X^{-p}$ in general.
A.2 Convexity of the objectives
In this subsection, we present the convexity proofs for all the considered objectives. Note that concavity of the $\log\det$ is well known and is stated here for completeness.
Lemma 9.
The function $f(X)=-\log\det(X)$ is convex for $X\in\mathbb{S}_{++}^{n}$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=-\log\det\left(A+tB\right)$.
$-\log\det(A+tB)$ $=-\log\det\left(A^{\frac{1}{2}}A^{\frac{1}{2}}+tA^{\frac{1}{2}}A^{-\frac{1}{2}}BA^{-\frac{1}{2}}A^{\frac{1}{2}}\right)$
$=-\log\det\left(A^{\frac{1}{2}}\left(I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}\right)A^{\frac{1}{2}}\right)$
$=-\log\det(A)-\log\det\left(I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}\right)$
Observe that $I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}$ is a positive definite matrix and its eigenvalues are $1+t\lambda_{i}$ with $\lambda_{i}$ being the $i$-th eigenvalue of $A^{-\frac{1}{2}}BA^{-\frac{1}{2}}$. Note that this also means $1+t\lambda_{i}>0$.
$=-\log\det(A)-\log\left(\prod\limits_{i=1}^{n}(1+t\lambda_{i})\right)$
$=-\log\det(A)-\sum\limits_{i=1}^{n}\log\left(1+t\lambda_{i}\right)$
The first and second derivatives of $g$ are
$h^{\prime}(t)$ $=-\sum\limits_{i=1}^{n}\frac{1}{1+t\lambda_{i}}\lambda_{i}$
$h^{\prime\prime}(t)$ $=\sum\limits_{i=1}^{n}\frac{\lambda_{i}^{2}}{\left(1+t\lambda_{i}\right)}$
$>0\quad\forall t$
Thus, $f(X)$ is convex. ∎
Lemma 10.
The function $g(X)=\operatorname{Tr}(X^{-p})$ is convex for $X\in\mathbb{S}_{++}^{n}$ and $p>0$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=\operatorname{Tr}\left(\left(A+tB\right)^{-p}\right)$.
$h(t)$ $=\operatorname{Tr}((A+tB)^{-p})$
$=\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}\right)$
Using Lemma˜7 for the derivative yields
$h^{\prime}(t)$ $=\operatorname{Tr}\left(p\left((A+tB)^{-1}\right)^{p-1}(-1)(A+tB)^{-1}B(A+tB)^{-1}\right)$
$=-p\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}\right)$
$h^{\prime\prime}(t)$ $=-p\operatorname{Tr}\left(p\left((A+tB)^{-1}\right)^{p-1}(-1)(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}\right$.
[…]
$=p\operatorname{Tr}\left((p+1)\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}B(A+tB)^{-1}\right)$
Restricting $h^{\prime\prime}$ to $t=0$ yields
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}\right).$
For the third derivative, we have
$h^{\prime\prime\prime}(t)$ $=p(p+1)\operatorname{Tr}\left(p\left((A+tB)^{-1}\right)^{p-1}(-1)(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}\right$.
[…]
$h^{\prime\prime\prime}(t)$ $=-p(p+1)(p+2)\operatorname{Tr}\left((A+tB)^{-p}B(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}\right).$
$h^{\prime\prime\prime}(t)_{\restriction t=0}$ $=-p(p+1)(p+2)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}BA^{-1}\right)$
By Corollary˜2, $A^{-p}$ is positive definite and thus, has a unique root $A^{-p/2}$. Using the traces cyclic property, we find
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}\right)$
$=p(p+1)\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}\right)$
$\geq 0$
The last inequality holds because the matrix $A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}$ is positive definite by positive definiteness of $A$, $B$ and $A^{-p}$. Thus, $g(X)$ is convex. ∎
Lemma 11.
The function $k(X)=\log\left(\operatorname{Tr}(X^{-p})\right)$ is convex on $\mathbb{S}_{++}^{n}$ for $p>0$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$.
$h(t)$ $=\log\left(\operatorname{Tr}\left((A+tB)^{-p}\right)\right)$
$h^{\prime}(t)$ $=-\frac{p}{\operatorname{Tr}((A+tB)^{-p})}\operatorname{Tr}((A+tB)^{-p-1}B)$
$h^{\prime\prime}(t)$ $=p\frac{(p+1)\operatorname{Tr}\left((A+tB)^{-p-2}BB\right)\operatorname{Tr}\left((A+tB)^{-p}\right)-p\operatorname{Tr}\left((A+tB)^{-p-1}B\right)^{2}}{\operatorname{Tr}\left((A+tB)^{-p}\right)^{2}}$
Restricting us to zero, we get
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p\frac{(p+1)\operatorname{Tr}\left(A^{-p-2}BB\right)\operatorname{Tr}\left(A^{-p}\right)-p\operatorname{Tr}\left(A^{-p-1}B\right)^{2}}{\operatorname{Tr}\left(A^{-p}\right)^{2}}.$
By assumption, $p$ is positive as is the denominator. For the nominator, we use the Cauchy Schwartz inequality $\operatorname{Tr}\left(XY^{\intercal}\right)^{2}\leq\operatorname{Tr}(XX^{\intercal})\operatorname{Tr}(YY^{\intercal})$ and set $X=A^{-p/2}$ and $Y=A^{-\frac{p+2}{2}}B$
$h^{\prime\prime}(t)_{\restriction t=0}$ $\geq 0$
Thus, $k(X)$ is convex. ∎
A.3 Additional smoothness proofs
In following, we show that all the considered functions are $L$-smooth if we can bound the minimum eigenvalue of the input matrix from below.
Lemma 12.
The function $f(X):=-\log\det(X)$ is $L$-smooth on $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\leq\lambda_{\min}(X)\}$ with $L\geq\frac{1}{\delta^{2}}$.
Proof.
Let $A\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma˜9, $f(X)$ is convex and the second derivative of $h$ is given by
$h^{\prime\prime}(t)$ $=\sum\limits_{i=1}^{n}\frac{\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}}{1+t\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)}$
It is sufficient to show that $h^{\prime\prime}(t)\restriction_{t=0}\leq L$ to argue smoothness for $f$.
$h^{\prime\prime}(t)\restriction_{t=0}$ $=\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}$
By using a similar idea to Lemma˜16, we find
$\leq\sum\limits_{i=1}^{n}\lambda_{\max}\left(A^{-1}\right)^{2}\lambda_{i}(B)^{2}$
$=\lambda_{\max}\left(A^{-1}\right)^{2}\operatorname{Tr}\left(B^{2}\right)$
$=\lambda_{\max}\left(A^{-1}\right)^{2}\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{1}{\lambda_{\min}(A)^{2}}$
$\geq\frac{1}{\delta^{2}}$
$=L$
Thus, the function $f$ is smooth at any point $A\in D$. ∎
Lemma 13.
For $p\in{\mathbb{R}}_{>0}$, the function $g(X):=\operatorname{Tr}\left(X^{-p}\right)$ is $L$-smooth on the domain $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\leq\lambda_{\min}(X)\}$ with
$L\geq\frac{p(p+1)}{\delta^{p+2}}.$
Proof.
Let $A\in D$ and $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. In Lemma˜10 the proof for convexity of $g(X)$ is presented. There one can also find the calculation for $h^{\prime\prime}(t)_{\restriction t=0}$.
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}\right).$
Since $A^{-p}$ is positive definite by Corollary˜2 and the trace is cyclic, we have
$=p(p+1)\operatorname{Tr}\left(A^{-\frac{p+2}{2}}BBA^{-\frac{p+2}{2}}\right)$
Note that $BB$ is postive semi-definite. Thus, we can use the bound in fang1994inequalities.
$\leq p(p+1)\lambda_{\max}\left(A^{-(p+2)}\right)\operatorname{Tr}(BB)$
$=p(p+1)\lambda_{\max}\left(A^{-(p+2)}\right)\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{p(p+1)}{\lambda_{\min}\left(A^{p+2}\right)}$
$\geq\frac{p(p+1)}{\delta^{p+2}}$
Thus, $g$ is smooth with $L=\frac{p(p+1)}{\delta^{p+2}}$. ∎
Lemma 14.
The function $k(X)=\log\left(\operatorname{Tr}(X^{-p})\right)$ is smooth on $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\alpha I\succcurlyeq X\succcurlyeq\delta I\}$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma˜11, we have
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p\frac{(p+1)\operatorname{Tr}\left(A^{-p-2}BB\right)\operatorname{Tr}\left(A^{-p}\right)-p\operatorname{Tr}\left(A^{-p-1}B\right)^{2}}{\operatorname{Tr}\left(A^{-p}\right)^{2}}$
$\leq p\frac{(p+1)\operatorname{Tr}\left(A^{-p}\right)\operatorname{Tr}\left(A^{-p-2}BB\right)}{\operatorname{Tr}(A^{-p})^{2}}$
$=p(p+1)\frac{\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-p-2}BB\right)}{\operatorname{Tr}\left(A^{-p}\right)}$
Note that $BB$ is PSD. The proof of Lemma˜16 can be adapted for one of matrices being symmetric and the other positive definite.
$\leq p(p+1)\frac{\lambda_{\max}\left(A^{-p-2}\right)\operatorname{Tr}(BB)}{\operatorname{Tr}\left(A^{-p}\right)}$
Observe that $Tr(BB)=\left\|B\right\|_{F}^{2}$. Without loss of generality, we can assume $\left\|B\right\|_{F}=1$.
$\leq p(p+1)\frac{\lambda_{\max}\left(A^{-p-2}\right)}{n\lambda_{\min}(A^{-p})}$
$=\frac{p(p+1)}{n}\frac{\lambda_{\max}(A^{p})}{\lambda_{\min}(A^{p+2})}$
$\leq\frac{p(p+1)}{n}\frac{\alpha^{p}}{\delta^{p+2}}$
$=L$
Remark.
In our case, the minimum eigenvalue is bounded around the optimum. However, it is not bounded in general for our feasible region unless we can introduce domain cuts. Hence, Lemmas˜14, 12 and 12 are not straight away useful to us.
A.4 Results on positive semi-definite matrices
We present two known results on the eigenvalues of the sum and product of positive definite matrices for completeness.
Lemma 15 (Courant-Fischer min-max Theorem).
Let $A,B\in{\mathbb{R}}^{n\times n}$ symmetric matrices and let $\lambda_{1}(A)\leq\dots\leq\lambda_{n}(A)$ and $\lambda_{1}(B)\leq\dots\leq\lambda_{n}(B)$ denote their respective eigenvalues in increasing order. Then, we have the following inequalities concerning the eigenvalues of their sum $A+B$:
$\lambda_{k}(A)+\lambda_{1}(B)\leq\lambda_{k}(A+B)\leq\lambda_{k}(A)+\lambda_{n}(B)$ (12)
Proof.
Firstly, we know that
$\lambda_{1}(A)$ $\leq\frac{\left\langle A,\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\leq\lambda_{n}(A)\quad\forall\mathbf{x}\in{\mathbb{R}}^{n}\backslash\{0\}.$
By Conrad-Fisher Min-Max Theorem, we have
$\lambda_{k}(A+B)$ $=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle(A+B)\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}+\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right).$
Most likely, there will not be an $\mathbf{x}$ attaining the maximum simultaneously for summands. Using the above inequality for the eigenvalues, we get
[…]
$=\lambda_{1}(B)+\lambda_{k}(A).$
The proof in the other direction follows analogously. ∎
Lemma 16.
Let $A,B\in\mathbb{S}_{++}^{n}$ and let $\lambda_{1}(A)\leq\dots\leq\lambda_{n}(A)$ and $\lambda_{1}(B)\leq\dots\leq\lambda_{n}(B)$ denote their respective eigenvalues in increasing order. Then, we have the following inequalities concerning the eigenvalues of $AB$:
$\lambda_{k}(A)\lambda_{1}(B)\leq\lambda_{k}(AB)\leq\lambda_{k}(A)\lambda_{n}(B)\quad\forall k\in[n].$ (13)
Proof.
Since $A$ and $B$ are both positive definite, they both have an invertible root, i.e. there exist matrices $\sqrt{A}$ and $\sqrt{B}$ with $A=\sqrt{A}\sqrt{A}$ and $B=\sqrt{B}\sqrt{B}$. Observe that $\lambda_{k}(AB)=\lambda_{k}(\sqrt{B}A\sqrt{B})$ and by definition of eigenvalues, we have
$\lambda_{1}(B)\leq\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\leq\lambda_{n}(B)\quad\forall\mathbf{x}\in{\mathbb{R}}^{n}\backslash\{0\}$
. Thus, we find
$\lambda_{k}(AB)$ $=\lambda_{k}(\sqrt{B}A\sqrt{B})$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle\sqrt{B}A\sqrt{B}\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\sqrt{B}\mathbf{x},\sqrt{B}\mathbf{x}\right\rangle}{\left\langle\sqrt{B}\mathbf{x},\sqrt{B}\mathbf{x}\right\rangle}\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
Using the above inequality for the eigenvalues of $B$ and the fact that the root of $B$ is invertible, we get
[…]
$=\lambda_{1}(B)\lambda_{k}(A)$
The other side of the statement follows analogously. Note that we technically only use that $B$ is regular. The matrix $A$ can be positive semi-definite. ∎
A.5 Notes on the Frobenius norm
Lemma 17.
Let $A\in{\mathbb{R}}^{m\times n}$ and $B\in{\mathbb{R}}^{n\times p}$. Then,
$\sigma_{\min}(A)\|B\|_{F}\leq\|AB\|_{F}\leq\sigma_{\max}\|B\|_{F}$ (14)
Proof.
By the definition of the Frobenius norm, we have
$\|AB\|_{F}^{2}$ $=\operatorname{Tr}\left((AB)^{\intercal}AB\right)$
$=\operatorname{Tr}(A^{\intercal}AB^{\intercal}B).$
Note that both $A^{\intercal}A$ and $B^{\intercal}B$ are positve semi-definite. Thus, we can use the result in fang1994inequalities:
$\geq\lambda_{\min}(A^{\intercal}A)\operatorname{Tr}(B^{\intercal}B)$
$=\sigma_{\min}(A)\|B\|_{F}^{2}.$
The norm is nonnegative in any case and by definition singular values are also nonnegative. Thus, we can take the root and get
$\sigma_{\min}(A)\|B\|_{F}\leq\|AB\|_{F}.$
The other direction follows analogously. ∎
B Custom branch-and-bound with coordinate descent for OEDP
In this section, we showcase the changes we made to the solver set-up in ahipacsaouglu2021branch. Firstly, we state and solve the problems of interest as minimization problems. Secondly, we forgo the $\mathbf{w}N\in{\mathbb{Z}}\,\forall i\in[m]$ in favor of directly having the variables $x$ to encode if and how often the experiments should be run. The D-Optimal Problem is then stated as
$\min_{\mathbf{x}}\,$ $\,-\log\det\left((X(\mathbf{x}))\right)$
s.t. $\sum\limits_{i=1}^{m}x_{i}=N$
$\mathbf{x}\geq 0$
$x_{i}\in{\mathbb{Z}}\,\forall i\in[m].$
The A-Optimal Problem is formulated as
$\min_{\mathbf{x}}\,$ $\,\log\left(\operatorname{Tr}(X^{-1})\right)$
s.t. $\sum\limits_{i=1}^{m}x_{i}=N$
$\mathbf{x}\geq 0$
$x_{i}\in{\mathbb{Z}}\,\forall i\in[m].$
Additionally, we have adapted the step size for the A-Optimal Problem. For the update in the objective, we get
$-\log(\operatorname{Tr}(X(\mathbf{x}_{+})))$ $=-\log(\operatorname{Tr}(X(\mathbf{x})^{-1})-\frac{\theta A+\theta^{2}B}{1+\theta C-\theta^{2}D})$
compare ahipacsaouglu2021branch. The constant $A$, $B$, $C$ and $D$ are defined as
$A$ $=\zeta_{j}-\zeta_{k}$
$B$ $=2\omega_{jk}\zeta_{jk}-\omega_{j}\zeta_{k}-\omega_{k}\zeta_{j}$
$C$ $=\omega_{j}-\omega_{k}$
$D$ $=\omega_{j}\omega_{k}-\omega_{jk}^{2}$
where $\omega_{j}=\mathbf{v}_{j}^{\intercal}X(x)^{-1}\mathbf{v}_{j}$ and $\zeta_{j}=\mathbf{v}_{j}^{\intercal}X(x)^{-2}\mathbf{v}_{j}$, and $\omega_{jk}=\mathbf{v}_{j}^{\intercal}X(x)^{-1}\mathbf{v}_{k}$ and $\zeta_{jk}=\mathbf{v}_{j}^{\intercal}X^{-2}\mathbf{v}_{k}$. The objective value should be improving.
$-\log\operatorname{Tr}(X(\mathbf{x})^{-1})$ $\geq-\log\operatorname{Tr}(X(\mathbf{x}_{+})^{-1})$
$\operatorname{Tr}(X(\mathbf{x})^{-1})$ $\leq\operatorname{Tr}(X(\mathbf{x}_{+})^{-1})$
$=\operatorname{Tr}(X(\mathbf{x})^{-1})-\frac{\theta A+\theta^{2}B}{1+\theta C-\theta^{2}D}$
Hence, we want to maximise the function $q(\theta)=\frac{\theta A+\theta^{2}B}{1+\theta C-\theta^{2}D}$. Let $\Delta=AD+BC$. By computing the first and second derivatives, it is easy to verify that we have
$\theta^{*}$ $=-\frac{B+\sqrt{B^{2}-A\Delta}}{\Delta}$
if $\Delta\neq 0$. If $\Delta=0$ but $B$ is not, then we have
$\theta^{*}$ $=-\frac{A}{2B}.$
If both $B$ and $\Delta$ are zero, then we choose $\theta$ as big as possible given the bound constraints.
$\theta$ $=\begin{cases}\min\left\{-\frac{B+\sqrt{B^{2}-A\Delta}}{\Delta},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta\neq 0\\ \min\left\{-\frac{A}{2B},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta=0\text{ and }B\neq 0\\ \min\{u_{j}-x_{j},x_{k}-l_{k}\}&\text{if }\Delta=B=0\text{ and }A>0\end{cases}$
C BPCG convergence under sharpness conditions
In this appendix, we prove linear convergence of BPCG on polytopes when the objective function is sharp or presents a Hölderian error bound. For the definition of sharpness, see Definition˜5. The following proofs rely on a geometric constant of a polytope $\mathcal{P}$.
Definition 7 (Pyramidal Witdh braun2022conditional).
The pyramidal width of a polytope $\mathcal{P}\subset{\mathbb{R}}^{n}$ is the largest number $\delta$ such that for any set $\mathcal{S}$ of some vertices of $\mathcal{P}$, we have
$\left\langle\mathbf{w},\mathbf{v}^{A}-\mathbf{v}^{FW}\right\rangle\geq\delta\frac{\left\langle\mathbf{w},\mathbf{x}-\mathbf{y}\right\rangle}{\left\|\mathbf{x}-\mathbf{y}\right\|}\quad\forall\mathbf{x}\in\operatorname{conv}\left(\mathcal{S}\right)\;\forall\mathbf{y}\in\mathcal{P}\;\forall\mathbf{w}\in{\mathbb{R}}^{n}$
where $\mathbf{v}^{A}:=\operatorname*{argmax}_{\mathbf{v}\in\mathcal{S}}\left\langle\mathbf{w},\mathbf{v}\right\rangle$ and $\mathbf{v}^{F}W:=\operatorname*{argmin}_{\mathbf{v}\in\mathcal{S}}\left\langle\mathbf{w},\mathbf{v}\right\rangle$.
Theorem 6 (Linear convergence of BPCG).
Let $P$ be a polytope of pyramidal width $\delta$ and diameter $D$, and let $f$ be an $L$-smooth and $(M,\theta)$-sharp function over $P$. Noting $f^{*}=\min_{\mathbf{x}\in P}f(\mathbf{x})$ and $\{\mathbf{x}_{t}\}_{i=0}^{T}$ the iterations of the BPCG algorithm, then it holds that:
$f(\mathbf{x}_{t})-f^{*}\leq\begin{cases}c_{0}(1-c_{1})^{2t-1}&\text{ if }1\leq t\leq\frac{t_{0}}{2}\\ \frac{(c_{1}/c_{2})^{1/\alpha}}{\left(1+c_{1}\alpha(2t-t_{0})\right)^{1/\alpha}}&\text{ if }t\geq\frac{t_{0}}{2}.\end{cases}$
with parameters:
$c_{1}=\frac{1}{2}$
$\alpha=2-2\theta$
$c_{2}=\frac{\delta^{2}}{8LD^{2}M^{2}}$
$t_{0}=\max\left\{\left\lfloor\mathrm{log}_{1-c_{1}}\left(\frac{(c_{1}/c_{2})^{1/\alpha}}{c_{0}}\right)\right\rfloor+2,1\right\}.$
Proof.
We will denote $h_{t}=f(\mathbf{x}_{t})-f^{*}$. The proof adapts that of tsuji2021sparser to the case of a sharp objective. tsuji2021sparser establishes that at any step $t$, the following holds:
$2\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle\geq\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{v}_{t}^{A}-\mathbf{v}_{t}^{FW}\right\rangle.$ (15)
Let us first assume that $t$ is not a drop step.
If $t$ follows Case (a) of Lemma˜19, then
$h_{t}-h_{t+1}$ $\geq\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle^{2}}{2LD^{2}}$ by (17)
$\geq\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{v}_{t}^{A}-\mathbf{v}^{FW}_{t}\right\rangle^{2}}{8LD^{2}}$ by (15)
[…]
If $t$ follows Case (b) of Lemma˜19, then:
$h_{t}-h_{t+1}$ $\geq\frac{1}{2}\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle\geq\frac{1}{2}h_{t},$
where the second inequality comes from $d_{t}$ being a FW step and from the FW gap being an upper bound on the primal gap.
Denoting with $\mathcal{T}_{\mathrm{drop}}\subseteq[T]$ the set of drop steps, we can already observe that the sequence $\{h_{t}\}_{t\in[T]\backslash\mathcal{T}_{\mathrm{drop}}}$ follows the hypotheses of Lemma˜20, with parameters $c_{1},c_{2},\alpha$ defined in the statement of the theorem.
We can then observe that the number of drop steps is bounded by the number of FW steps on the sequence. This implies in the worst case that in $2t,t>1$ iterations, we can reach the primal gap of $t$ FW and pairwise steps combined. Let us denote $\tau$ the number of steps that were not drop steps up to iteration $t$, then:
$h_{t}$ $\leq\begin{cases}c_{0}(1-c_{1})^{\tau-1}&1\leq\tau\leq t_{0}\\ \frac{(c_{1}/c_{2})^{1/\alpha}}{\left(1+c_{1}\alpha(\tau-t_{0})\right)^{1/\alpha}}&\tau\geq t_{0},\end{cases}$
$\leq\begin{cases}c_{0}(1-c_{1})^{2t-1}&1\leq t\leq\frac{t_{0}}{2}\\ \frac{(c_{1}/c_{2})^{1/\alpha}}{\left(1+c_{1}\alpha(2t-t_{0})\right)^{1/\alpha}}&t\geq\frac{t_{0}}{2}.\end{cases}$
with $t_{0}$ as defined in (19).
Lemma 18 (Geometric sharpness braun2022conditional, Lemma 3.32).
Let $P$ be a polytope with pyramidal width $\delta$ and let $f$ be $(M,\theta)$-sharp over $P$. Given $S\subseteq\mathrm{Vert}(P)$ such that $\mathbf{x}\in\mathrm{conv}(S)$ and:
$\mathbf{v}^{FW}\in\operatorname*{argmin}_{\mathbf{v}\in P}\left\langle\nabla f(\mathbf{x}),\mathbf{v}\right\rangle$
$\mathbf{v}^{A}\in\operatorname*{argmax}_{\mathbf{v}\in P}\left\langle\nabla f(\mathbf{x}),\mathbf{v}\right\rangle.$
Then:
$\frac{\delta}{M}(f(\mathbf{x})-f^{*})^{(1-\theta)}\leq\left\langle\nabla f(\mathbf{x}),\mathbf{v}^{A}-\mathbf{v}^{FW}\right\rangle.$ (16)
Lemma 19 (Minimum decrease tsuji2021sparser, Lemma 3.4).
Suppose that $t$ is not a drop step, and let $\alpha_{t}^{*}=\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle}{L\left\|\mathbf{d}_{t}\right\|^{2}}$.
  1. (a)
    If step $t$ is either a FW step with $\alpha_{t}^{*}<1$ or a pairwise step, then:
    $f(\mathbf{x}_{t})-f(\mathbf{x}_{t+1})\geq\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle^{2}}{2LD^{2}}.$ (17)
  2. (b)
    If step $t$ is a FW step with $\alpha_{t}>1$, we have:
    $f(\mathbf{x}_{t})-f(\mathbf{x}_{t+1})\geq\frac{1}{2}\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle.$ (18)
Lemma 20 (Convergence rates from contraction inequalities, braun2022conditional, Lemma 2.21).
Let $\{h_{t}\}_{t\geq 1}$ be a sequence of positive numbers and $c_{0},c_{1},c_{2},\alpha$ such that:
$\begin{cases}c_{1}<1\\ h_{1}\leq c_{0}\\ h_{t}-h_{t+1}\geq h_{t}\min\{c_{1},c_{2}h_{t}^{\alpha}\}\,\forall t\geq 1,\end{cases}$
then,
$h_{t}\leq\begin{cases}c_{0}(1-c_{1})^{t-1}&1\leq t\leq t_{0}\\ \frac{(c_{1}/c_{2})^{1/\alpha}}{\left(1+c_{1}\alpha(t-t_{0})\right)^{1/\alpha}}&t\geq t_{0},\end{cases}$
where
$t_{0}=\max\left\{\left\lfloor\mathrm{log}_{1-c_{1}}\left(\frac{(c_{1}/c_{2})^{1/\alpha}}{c_{0}}\right)\right\rfloor+2,1\right\}.$ (19)
Finally, we can exploit the recent results of zhao2025new on AFW for log-homogeneous self-concordant barrier functions to show that on D-Opt Problem, BPCG converges at a linear rate in primal and FW gaps.
Proposition 1 (Linear convergence of BPCG on D-Opt (zhao2025new)).
The blended pairwise conditional gradient contracts the primal gap $h_{t}$ at a linear rate $h_{t}\leq h_{0}\exp(-ct)$ with $h_{0}$ the primal gap of the initial domain-feasible solution and $c$ a constant depending on the design matrix and polytope.
Proof.
The core observation required to extend zhao2025new is that the gap produced by BPCG upper-bounds the one produced by AFW. Let $\mathbf{d}^{\text{BPCG}},\mathbf{d}^{\text{AFW}}$ be the search directions produced by BPCG, AFW respectively. A drop step can occur in both algorithms in at most half of the iterations, since each drop step reduces the size of the active set by one. We can focus on “effective” steps in which no vertex is dropped. BPCG either takes a local pairwise step with direction:
$\mathbf{d}^{\text{BPCG}}=\mathbf{a}-\mathbf{s}$
with $\mathbf{a}=\operatorname*{argmax}_{\mathbf{v}\in\mathcal{S}_{t}}\left\langle\nabla f(\mathbf{x}),\mathbf{v}\right\rangle$
$\mathbf{s}=\operatorname*{argmin}_{\mathbf{v}\in\mathcal{S}_{t}}\left\langle\nabla f(\mathbf{x}),\mathbf{v}\right\rangle,$
and $\mathcal{S}_{t}$ the local active set. Using the same notation, $\mathbf{d}^{\text{AFW}}=\mathbf{a}-\mathbf{x}$. The quantity determining whether a local pairwise step is selected is $\left\langle\nabla f(\mathbf{x}),\mathbf{d}^{\text{BPCG}}\right\rangle$, which needs to be greater than the FW gap $\left\langle\nabla f(\mathbf{x}),\mathbf{x}-\mathbf{v}\right\rangle$ at the current point. Regardless of the direction taken by BPCG, tsuji2021sparser provides:
$\left\langle\nabla f(\mathbf{x}),\mathbf{d}\right\rangle\geq 1/2\left\langle\nabla f(\mathbf{x}),\mathbf{a}-\mathbf{v}\right\rangle$
which is precisely the only inequality required from effective AFW steps. The rest of the proofs from zhao2025new follow without relying on the AFW descent direction specifically. ∎
Proposition˜1 unfortunately relies on the fact that the D-Opt objective is a log-barrier in addition to being self-concordant. Since this is not the case for the trace inverse function, we cannot transpose this result to the A-Opt Problem.
D Computational experiments
In this appendix section, we present additional computational experiments. Table˜5 show an extended comparison of the performance between the different solution approaches on the A-Optimal Problem, D-Optimal Problem, A-Fusion Problem and D-Fusion Problem. In Tables˜6 and 7, the results of the computations on the GTI-Optimal and GTI-Fusion Problem are shown. The Tables˜8, 9 and 10 compare the performance of Boscia using different settings.
Table 5: Comparing Boscia, SCIP, Co-BnB and the direct conic formulation and second-order formulation using Pajarito on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table.
The average of the relative gap does not include instances that were solved by all methods. The average of number nodes/cuts is only taken for the instances solved in that group. The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
Boscia Co-BnB Direct Conic SOCP SCIP OA
Type Corr. Solved after (s) % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # cuts % solved Time (s) relative gap # cuts % solved Time (s) relative gap # cuts
AO no 0 52 % 374.74 0.0851 7340 48 % 401.67 51.4694 22211 8 % 2382.52 2.4436 112 20 % 1578.03 Inf 2842 - - - -
10 40 % 1230.45 0.0964 11775 35 % 1309.24 56.5936 37906 0 % 3600.09 3.3563 - 0 % 3600.02 Inf - - - - -
100 25 % 2974.04 0.118 22241 19 % 2922.12 66.0235 81489 0 % 3600.11 0.6943 - 0 % 3600.01 Inf - - - - -
1000 17 % 3164.43 0.1291 19766 10 % 3360.8 71.1272 86515 0 % 3600.12 0.7501 - 0 % 3600.01 Inf - - - - -
2000 4 % 3583.85 0.1482 22096 0 % 3602.18 74.692 - 0 % 3600.09 0.8378 - 0 % 3600.01 Inf - - - - -
AO yes 0 68 % 227.69 0.0336 7814 60 % 301.8 2058.4159 22379 20 % 1515.23 30.3669 188 20 % 1861.35 Inf 3623 - - - -
10 60 % 481.42 0.0348 10181 50 % 864.83 2058.4159 33031 2 % 3149.17 34.3286 206 5 % 3368.07 Inf 5752 - - - -
100 45 % 1421.72 0.0434 16675 31 % 2476.41 2058.4159 59473 0 % 3600.05 0.4903 - 0 % 3600.01 Inf - - - - -
1000 16 % 3544.51 0.061 49147 5 % 3583.18 2107.5216 89805 0 % 3600.05 1.2972 - 0 % 3600.01 Inf - - - - -
AF no 0 80 % 62.83 0.0168 11634 80 % 50.92 30.35 10648 4 % 2846.5 0.0371 58 26 % 1686.2 Inf 5634 38 % 463.65 0.0302 801
10 66 % 460.44 0.0213 22887 66 % 409.38 37.6051 21718 0 % 3600.03 0.0548 - 0 % 3600.09 Inf - 7 % 3172.81 0.0379 1930
100 47 % 1644.98 0.0272 41058 47 % 1374.15 42.7763 33254 0 % 3600.04 0.074 - 0 % 3600.14 Inf - 0 % 3600.1 0.0485 -
1000 25 % 3087.98 0.0372 67544 17 % 3125.27 48.0929 67087 0 % 3600.06 0.0967 - 0 % 3600.17 Inf - 0 % 3600.15 0.0646 -
2000 11 % 3416.18 0.0461 70407 0 % 3601.84 51.6913 - 0 % 3600.06 0.1108 - 0 % 3600.22 Inf - 0 % 3600.19 0.071 -
AF yes 0 26 % 1368.83 2.1573 28578 52 % 311.36 1407.4062 30449 24 % 1148.87 9.0208 318 20 % 2131.81 Inf 4539 14 % 1480.73 3.4254 552
10 11 % 3030.17 2.5452 51328 37 % 1336.22 1468.0559 56106 3 % 3418.49 10.6718 1086 0 % 3603.12 Inf - 0 % 3600.04 3.6558 -
100 6 % 3315.04 2.8901 58354 27 % 2303.15 1468.0559 84247 0 % 3600.08 12.2264 - 0 % 3603.59 Inf - 0 % 3600.05 3.9173 -
1000 0 % 3600.06 3.2127 - 17 % 3079.19 1468.0559 128319 0 % 3600.09 13.8081 - 0 % 3604.08 Inf - 0 % 3600.05 4.2154 -
2000 0 % 3600.07 3.6098 - 4 % 3595.22 1533.847 154005 0 % 3600.09 13.7875 - 0 % 3604.73 Inf - 0 % 3600.06 4.5519 -
DF no 0 94 % 3.49 0.0102 301 86 % 22.79 0.0382 6782 12 % 1906.61 0.0126 433 18 % 2213.03 Inf 4787 50 % 324.11 0.0133 1155
10 50 % 727.47 0.0128 1424 17 % 2540.85 0.1531 32841 0 % 3879.62 0.0201 - 0 % 3600.0 Inf - 0 % 3600.17 0.0163 -
100 40 % 1176.39 0.0133 94 0 % 3602.41 0.1825 - 0 % 3936.34 0.0213 - 0 % 3600.0 Inf - 0 % 3600.2 0.0176 -
DF yes 0 60 % 54.01 0.0842 7635 56 % 153.26 0.4948 22485 6 % 2284.06 0.1556 319 6 % 2934.61 Inf 3117 28 % 752.15 0.092 786
10 17 % 2634.14 0.1592 56466 12 % 2705.02 0.948 162320 0 % 3041.2 0.302 - 0 % 3600.0 Inf - 0 % 3600.05 0.1679 -
100 14 % 3251.87 0.161 74805 9 % 3265.96 0.9635 239147 0 % 3603.96 0.258 - 0 % 3600.0 Inf - 0 % 3600.05 0.1711 -
1000 5 % 3535.72 0.1761 40687 0 % 3599.62 1.0588 - 0 % 3604.36 0.2811 - 0 % 3600.0 Inf - 0 % 3600.05 0.185 -
DO no 0 68 % 136.92 0.0208 3578 58 % 216.31 2.7678 9559 20 % 1241.37 0.676 1077 20 % 2169.4 Inf 4514 - - - -
10 56 % 618.39 0.0226 5916 42 % 991.36 3.3814 17925 3 % 2355.99 0.8846 3056 0 % 3600.0 Inf - - - - -
100 45 % 1330.76 0.0257 8559 28 % 1978.6 4.5636 27449 3 % 2126.78 1.3081 3056 0 % 3600.0 Inf - - - - -
1000 17 % 3261.11 0.0358 18628 0 % 3602.01 7.4215 - 0 % 2741.23 4.4622 - 0 % 3600.01 Inf - - - - -
2000 12 % 3515.21 0.0294 30726 0 % 3601.85 3.4692 - 0 % 3614.18 4.4622 - 0 % 3600.01 Inf - - - - -
DO yes 0 100 % 2.56 0.0094 108 70 % 101.89 0.0809 7973 10 % 1398.01 0.0206 309 10 % 2526.74 Inf 2871 - - - -
10 100 % 65.51 0.0099 439 0 % 3599.94 0.2101 - 0 % 1347.23 Inf - 0 % 3600.01 Inf - - - - -
100 100 % 195.36 0.0099 828 0 % 3600.17 0.1967 - 0 % 1204.26 Inf - 0 % 3600.02 Inf - - - - -
Table 6: Comparing Boscia and Co-BnB on the General-Trace-Inverse (GTI) Optimal Problem for different values of $p$ and under independent and correlated data. Note the the objective function has the form $\phi_{p}(X(\mathbf{x}))=-1/p\log\left(\operatorname{Tr}\left(X(\mathbf{x})^{p}\right)\right)$.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table.
The relative gap is computed for the instances on which at least one method did not terminate within the time limit. That means it excludes the instances on which all methods terminated. The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
Boscia Boscia Log Co-BnB
p Corr. Solved after (s) # inst. % solved Time (s) % solved Time (s) % solved Time (s)
GTI_25 no 0 30 67 % 162.41 87 % 27.13 50 % 359.92
GTI_25 no 10 16 38 % 1880.31 75 % 200.69 12 % 2864.51
GTI_25 no 100 9 11 % 3437.77 56 % 930.74 0 % 3622.39
GTI_25 no 1000 5 0 % 3606.39 20 % 3222.96 0 % 3640.97
GTI_25 yes 0 30 67 % 110.25 77 % 85.83 60 % 264.22
GTI_25 yes 10 16 38 % 1461.4 56 % 896.41 25 % 2614.2
GTI_25 yes 100 14 29 % 2309.53 50 % 1437.34 14 % 3307.62
GTI_25 yes 1000 7 0 % 3600.83 0 % 3600.9 0 % 3599.3
GTI_50 no 0 30 53 % 246.61 70 % 85.87 50 % 430.08
GTI_50 no 10 17 18 % 3139.96 47 % 1197.73 12 % 3456.55
GTI_50 no 100 14 0 % 3601.04 36 % 2242.33 0 % 3611.02
GTI_50 no 1000 10 0 % 3601.45 10 % 3581.72 0 % 3615.73
GTI_50 yes 0 30 60 % 129.86 57 % 245.36 57 % 313.7
GTI_50 yes 10 21 43 % 675.52 38 % 1340.71 38 % 1326.2
GTI_50 yes 100 16 25 % 2105.9 19 % 3238.0 19 % 3193.81
GTI_75 no 0 30 37 % 799.73 43 % 505.86 50 % 471.95
GTI_75 no 10 25 24 % 2171.71 32 % 1308.01 40 % 1228.42
GTI_75 no 100 21 10 % 3011.64 19 % 2210.64 29 % 2093.83
GTI_75 no 1000 17 0 % 3601.73 0 % 3602.06 12 % 3518.93
GTI_75 yes 0 30 47 % 541.67 43 % 557.15 50 % 357.77
GTI_75 yes 10 25 36 % 1514.11 32 % 1626.43 40 % 904.04
GTI_75 yes 100 21 24 % 2446.72 19 % 2739.52 29 % 1608.59
GTI_75 yes 1000 16 0 % 3600.46 0 % 3600.45 6 % 3440.58
GTI_75 yes 2000 15 0 % 3600.48 0 % 3600.48 0 % 3599.49
GTI_150 no 0 30 43 % 659.25 40 % 576.51 50 % 619.95
GTI_150 no 10 25 32 % 1908.22 28 % 1673.37 40 % 1669.28
GTI_150 no 1000 17 0 % 3600.54 0 % 3601.19 12 % 3571.66
GTI_150 yes 0 30 47 % 418.43 53 % 306.7 47 % 506.3
GTI_150 yes 10 24 33 % 1413.27 42 % 1076.7 33 % 1539.25
GTI_150 yes 100 19 16 % 2602.93 26 % 2474.23 16 % 2546.07
GTI_150 yes 1000 16 0 % 3600.04 12 % 3520.88 0 % 3599.27
GTI_200 no 0 30 43 % 661.29 50 % 562.4 50 % 630.18
GTI_200 no 10 25 32 % 1880.56 40 % 1601.07 40 % 1642.44
GTI_200 no 100 25 32 % 1880.56 40 % 1601.07 40 % 1642.44
GTI_200 no 1000 18 6 % 3510.95 17 % 3342.37 17 % 3207.77
GTI_200 no 2000 16 0 % 3601.01 6 % 3590.35 6 % 3521.64
GTI_200 yes 0 30 43 % 644.53 53 % 297.98 47 % 611.8
GTI_200 yes 10 24 29 % 2068.4 42 % 1062.97 33 % 1851.58
GTI_200 yes 100 20 20 % 2632.63 30 % 2026.91 20 % 2734.27
GTI_200 yes 1000 16 0 % 3600.06 12 % 3547.5 0 % 3599.23
Table 7: Comparing Boscia, Co-BnB and SCIP on the General-Trace-Inverse (GTI) Fusion Problem for different values of $p$ and under independent and correlated data. Note the the objective function has the form $\phi_{p}(X(\mathbf{x}))=-1/p\log\left(\operatorname{Tr}\left(C+X(\mathbf{x})^{p}\right)\right)$.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table.
The relative gap is computed for the instances on which at least one method did not terminate within the time limit. That means it excludes the instances on which all methods terminated. The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
Boscia Boscia Log Co-BnB SCIP OA
p Corr. Solved after (s) # inst. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
GTIF_25 no 0 30 90 % 6.57 100 % 1.71 50 % 725.52 47 % 376.08
GTIF_25 no 10 3 67 % 470.21 100 % 56.43 0 % 3755.69 0 % 3600.18
GTIF_25 no 100 1 0 % 3602.77 100 % 137.03 0 % 3834.75 0 % 3600.01
GTIF_25 yes 0 30 53 % 138.78 57 % 184.83 27 % 1449.73 27 % 927.92
GTIF_25 yes 10 19 26 % 1299.73 32 % 1260.16 0 % 3601.55 0 % 3600.03
GTIF_25 yes 100 16 12 % 2722.54 19 % 2572.48 0 % 3601.84 0 % 3600.04
GTIF_25 yes 1000 14 0 % 3600.55 7 % 3566.88 0 % 3602.03 0 % 3600.04
GTIF_50 no 0 30 87 % 9.1 90 % 3.63 47 % 787.45 43 % 349.14
GTIF_50 no 10 6 33 % 1865.24 50 % 310.51 0 % 3885.14 0 % 3600.18
GTIF_50 yes 0 30 57 % 175.85 53 % 271.12 30 % 1547.49 20 % 850.41
GTIF_50 yes 10 21 38 % 1010.86 33 % 1617.58 0 % 3601.95 0 % 3600.03
GTIF_50 yes 100 16 19 % 2753.86 12 % 3261.97 0 % 3602.31 0 % 3600.04
GTIF_50 yes 1000 14 7 % 3489.55 7 % 3569.66 0 % 3602.47 0 % 3600.04
GTIF_75 no 0 30 70 % 85.39 87 % 42.47 57 % 353.3 43 % 415.34
GTIF_75 no 10 15 60 % 362.32 73 % 144.1 33 % 1347.24 27 % 1927.57
GTIF_75 no 100 6 17 % 2549.51 33 % 2006.07 0 % 3832.95 0 % 3600.04
GTIF_75 no 1000 5 0 % 3603.61 20 % 3333.88 0 % 3881.18 0 % 3600.03
GTIF_75 yes 0 30 37 % 741.57 43 % 434.78 27 % 1639.63 23 % 1053.8
GTIF_75 yes 10 24 21 % 2351.14 29 % 1512.43 12 % 3038.52 8 % 2837.69
GTIF_75 yes 100 20 10 % 3212.49 15 % 2855.63 0 % 3602.37 0 % 3600.05
GTIF_75 yes 1000 18 6 % 3567.81 6 % 3373.72 0 % 3602.86 0 % 3600.06
GTIF_150 no 0 30 77 % 91.83 67 % 129.57 47 % 780.56 40 % 406.94
GTIF_150 no 10 19 63 % 629.52 47 % 821.17 21 % 2030.26 11 % 2732.57
GTIF_150 no 100 14 50 % 1556.16 36 % 1922.49 0 % 3712.95 0 % 3600.09
GTIF_150 no 1000 9 22 % 3112.4 11 % 3171.4 0 % 3776.35 0 % 3600.13
GTIF_150 no 2000 8 12 % 3488.09 0 % 3601.58 0 % 3798.53 0 % 3600.14
GTIF_150 yes 0 30 37 % 715.74 50 % 214.97 30 % 1415.97 23 % 885.83
GTIF_150 yes 10 20 15 % 2749.79 25 % 1837.42 5 % 2727.35 0 % 3600.04
GTIF_150 yes 100 17 18 % 2622.1 18 % 2813.97 0 % 3603.41 0 % 3600.04
GTIF_150 yes 1000 15 7 % 3379.09 7 % 3515.28 0 % 3603.8 0 % 3600.05
GTIF_200 no 0 30 73 % 96.28 70 % 117.1 57 % 797.56 40 % 402.28
GTIF_200 no 10 20 60 % 554.66 55 % 677.02 35 % 2413.83 10 % 2710.21
GTIF_200 no 100 15 47 % 1536.15 40 % 1840.11 13 % 3313.81 0 % 3600.06
GTIF_200 no 1000 9 11 % 3319.15 0 % 3601.02 0 % 3762.3 0 % 3600.09
GTIF_200 yes 0 30 33 % 905.59 53 % 202.0 33 % 1305.28 20 % 923.45
GTIF_200 yes 10 19 21 % 2561.54 32 % 1658.86 5 % 3586.45 0 % 3600.05
GTIF_200 yes 100 18 22 % 2513.56 28 % 2167.04 6 % 3585.71 0 % 3600.06
GTIF_200 yes 1000 14 7 % 3471.17 7 % 3513.18 0 % 3600.85 0 % 3600.07
Table 8: Comparing the different Boscia settings on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data.
The settings displayed here are the default, so all options are off, using heuristics, using SCIP to model the constraints and tightening the lazifaction for the BPCG algorithm solving the nodes.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table. A line will also be omitted if nothing changes compared to the previous one.
The relative gap only includes the instances on which at least one run didn’t terminate within the time limit. That means it does not include the instances that were solved by all.
The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
Default Heuristics Shadow Set MIP SCIP
Type Corr. Solved after (s) # inst. % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes
AO no 0 18 50 % 358.32 7.5254 7775 50 % 363.71 0.1521 8291 50 % 365.61 5.2159 8257 44 % 560.05 5.2295 2368
10 13 31 % 1410.24 10.4161 14871 31 % 1450.37 0.2069 16194 31 % 1450.66 7.2184 16044 23 % 1820.8 7.2372 2970
100 10 10 % 3554.77 13.5379 50135 10 % 3587.62 0.266 55115 10 % 3594.98 9.3809 54589 0 % 3600.25 9.4054 -
AO yes 0 18 44 % 831.95 0.1441 39425 44 % 778.14 0.2456 33389 44 % 823.4 0.2557 35199 44 % 1434.99 0.3364 34942
10 12 17 % 3066.58 0.2783 93409 17 % 2994.35 0.3634 84460 17 % 2944.02 0.3785 71701 17 % 3452.15 0.4996 79349
AF no 0 18 83 % 64.13 0.0196 15006 83 % 66.21 0.018 14433 83 % 64.07 0.017 14059 78 % 175.18 0.0219 10679
10 11 73 % 446.0 0.0259 27310 73 % 440.79 0.0232 26154 73 % 427.53 0.0216 25479 64 % 1023.65 0.0296 20399
100 8 62 % 1216.92 0.0318 40277 62 % 1179.39 0.0281 38299 62 % 1126.97 0.0259 37212 50 % 2145.02 0.0369 31333
1000 4 25 % 3097.25 0.0537 64107 25 % 2905.95 0.0463 59519 25 % 3012.72 0.0418 62081 0 % 3600.69 0.0639 -
AF yes 0 18 28 % 1330.47 2.1036 31537 28 % 1386.95 2.1095 32697 28 % 1369.5 2.0899 32533 28 % 1919.67 2.3779 27555
10 17 24 % 1847.36 2.2267 39036 24 % 1933.2 2.233 40487 24 % 1902.13 2.2122 40282 24 % 2481.32 2.5172 34050
100 15 13 % 3107.54 2.5223 63405 13 % 3059.65 2.5294 58879 13 % 3060.47 2.5059 60458 13 % 3431.17 2.8515 51914
1000 14 7 % 3441.35 2.7017 87165 7 % 3370.08 2.7094 75443 7 % 3399.93 2.6842 80221 7 % 3593.22 3.0545 69063
DF no 0 18 94 % 4.76 0.0101 219 94 % 4.54 0.0098 217 94 % 4.43 0.0102 232 94 % 11.15 0.0103 252
10 3 67 % 575.13 0.0135 125 67 % 419.91 0.0119 97 67 % 356.85 0.013 89 67 % 778.56 0.0151 118
DF yes 0 18 61 % 77.57 0.097 18067 61 % 78.71 0.097 18990 61 % 79.39 0.0968 18161 50 % 105.04 0.1012 169
10 10 30 % 1942.63 0.1673 66153 30 % 2001.29 0.1671 69540 30 % 1940.96 0.1669 66478 10 % 2372.28 0.1749 1243
100 9 22 % 3323.61 0.1848 98596 22 % 3437.13 0.1845 103673 22 % 3320.6 0.1844 99082 0 % 3600.02 0.1932 -
2000 8 12 % 3455.16 0.2067 161153 12 % 3452.91 0.2064 161139 12 % 3454.46 0.2062 161069 0 % 3600.02 0.2161 -
DO no 0 18 67 % 195.53 0.024 11231 67 % 187.77 0.0238 11117 67 % 188.2 0.0245 11062 56 % 400.29 0.0246 3751
10 11 45 % 1468.68 0.0331 25307 45 % 1394.14 0.0327 25024 45 % 1449.37 0.0338 24941 27 % 2057.64 0.0341 9819
100 8 25 % 3389.76 0.0418 48615 25 % 3309.23 0.0413 48268 25 % 3447.06 0.0428 48411 0 % 3600.31 0.0431 -
1000 6 17 % 3549.76 0.0495 54787 17 % 3470.1 0.0483 54409 17 % 3597.51 0.0504 55215 0 % 3600.4 0.0506 -
2000 5 20 % 3539.8 0.0552 54787 20 % 3444.68 0.0539 54409 20 % 3597.01 0.0567 55215 0 % 3600.47 0.0564 -
DO yes 0 18 100 % 5.47 0.0095 494 100 % 5.05 0.0095 563 100 % 4.83 0.0094 385 100 % 9.75 0.0095 436
10 4 100 % 230.32 0.01 1980 100 % 170.88 0.0098 2232 100 % 149.39 0.01 1502 100 % 333.38 0.0099 1720
Table 9: Comparing the different Boscia settings on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data.
The settings displayed here are using tightening of the bounds, enabling the shadow set within BPCG and using Strong Branching instead of the most fractional branching rule.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table. A line will also be omitted if nothing changes compared to the previous one.
The relative gap only includes the instances on which at least one run didn’t terminate within the time limit. That means it does not include the instances that were solved by all.
The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
Strong Branching Tightening Tighten Lazification Secant Blended Conditional Gradient
Type Corr. Solved after (s) # inst. % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes % solved Time (s) relative gap # nodes
AO no 0 18 44 % 509.39 7.7803 2170 50 % 353.06 7.5423 7847 50 % 335.12 7.562 8113 44 % 395.39 7.5383 2369 50 % 288.11 0.7865 8233
10 13 23 % 1785.5 10.7691 2848 31 % 1399.5 10.4396 15068 31 % 1327.05 10.4668 15506 23 % 1498.71 10.434 2891 31 % 1156.73 1.0854 15548
100 10 0 % 3635.44 13.9969 - 10 % 3563.21 13.5685 51487 10 % 3469.34 13.6038 52459 0 % 3600.14 13.5612 - 10 % 3226.67 1.408 51685
AO yes 0 18 33 % 1411.93 0.5182 26947 44 % 895.28 0.2547 35056 44 % 1032.46 0.3766 52251 44 % 780.03 0.255 44423 39 % 1056.59 0.2829 50528
10 12 8 % 3482.79 0.9416 69367 17 % 2859.77 0.377 66564 17 % 3262.57 0.5599 109156 17 % 3106.83 0.3776 98795 8 % 3475.97 0.5559 135093
AF no 0 18 72 % 224.3 0.0402 4937 83 % 69.05 0.0184 15279 78 % 77.05 0.0196 16534 78 % 88.49 0.0198 12429 83 % 74.95 0.0191 23414
10 11 55 % 1382.75 0.0595 9816 73 % 471.19 0.0239 27754 64 % 563.16 0.0258 31716 64 % 673.09 0.0261 23944 73 % 513.29 0.0249 42680
100 8 38 % 2711.25 0.0781 13742 62 % 1281.81 0.0291 41038 50 % 1540.09 0.0317 49409 50 % 1805.8 0.0322 37498 62 % 1471.05 0.0305 63486
1000 4 0 % 3645.98 0.1458 - 25 % 3011.79 0.0485 60591 25 % 3277.73 0.0533 94439 0 % 3600.18 0.0544 - 25 % 3464.55 0.051 119747
AF yes 0 18 28 % 1658.47 3.2623 30096 28 % 1325.29 2.1184 28058 28 % 1461.5 2.4675 36446 28 % 1296.61 2.1909 30838 28 % 1378.94 2.1662 35027
10 17 24 % 2198.67 3.4656 37295 24 % 1824.24 2.2424 34688 24 % 2016.49 2.6121 45073 24 % 1855.8 2.3194 38217 24 % 1899.87 2.3009 43212
100 15 13 % 3420.23 3.9592 65163 13 % 3000.06 2.54 55034 13 % 3138.87 2.9591 68958 13 % 3062.15 2.6273 56098 13 % 2985.76 2.6282 62897
1000 14 7 % 3499.51 4.263 75549 7 % 3370.29 2.7208 73985 7 % 3434.08 3.1697 87381 7 % 3412.79 2.8142 79157 7 % 3332.5 2.8296 78621
DF no 0 18 89 % 25.69 0.0125 272 94 % 4.94 0.0104 227 94 % 5.24 0.0102 333 94 % 5.27 0.0105 209 94 % 5.61 0.0103 331
10 3 33 % 3300.66 0.0281 79 67 % 447.07 0.0152 101 67 % 614.7 0.0138 233 67 % 451.16 0.0142 99 67 % 739.18 0.0126 232
DF yes 0 18 56 % 133.22 0.1285 16635 61 % 77.67 0.0967 18104 61 % 76.96 0.0962 18141 56 % 85.74 0.0987 16260 61 % 72.59 0.0938 17502
10 10 20 % 2517.61 0.2244 83020 30 % 1935.53 0.1668 66290 30 % 1838.5 0.1656 66376 20 % 2133.82 0.1703 81149 30 % 1648.16 0.1615 64030
100 9 11 % 3493.0 0.2482 163999 22 % 3323.78 0.1842 98814 22 % 3215.14 0.1829 98904 11 % 3495.7 0.1882 161073 22 % 2884.32 0.1784 95276
2000 8 12 % 3479.82 0.2777 163999 12 % 3460.11 0.206 161077 12 % 3454.03 0.2045 161061 12 % 3482.87 0.2104 161073 12 % 3362.42 0.1995 161107
DO no 0 18 56 % 354.36 0.1266 3254 61 % 196.4 0.0243 7423 67 % 191.89 0.0248 13438 61 % 227.48 0.0252 7050 67 % 158.11 0.0221 14300
10 11 27 % 2133.07 0.2009 8332 36 % 1503.71 0.0336 18386 45 % 1429.48 0.0343 30310 36 % 1652.29 0.0354 17358 45 % 1156.8 0.03 32303
100 8 0 % 3659.93 0.2725 - 12 % 3473.32 0.0425 43491 25 % 3393.4 0.0434 58562 12 % 3577.58 0.045 42633 25 % 3006.11 0.0375 63716
1000 6 0 % 3679.97 0.3538 - 0 % 3601.58 0.0499 - 17 % 3540.67 0.0516 65771 0 % 3600.1 0.0535 - 17 % 3315.64 0.0442 74469
2000 5 0 % 3696.16 0.4197 - 0 % 3601.9 0.0561 - 20 % 3528.91 0.058 65771 0 % 3600.12 0.0599 - 20 % 3261.52 0.0499 74469
DO yes 0 18 94 % 18.54 0.0094 281 100 % 5.63 0.0093 378 100 % 5.5 0.0097 638 100 % 5.67 0.0094 392 94 % 4.76 0.0095 280
10 4 75 % 1385.8 0.0104 1206 100 % 205.16 0.0098 1456 100 % 204.36 0.01 2491 100 % 221.99 0.0098 1470 75 % 134.78 0.0098 1108
Table 10: Comparing performance between A-criterion with and without $\log$. For $\operatorname{Tr}\left(X^{-1}\right)$, we use both the Secant linesearch as well as the Adaptive linesearch wrapped in a MonotonicGeneric linesearch. For $\log\operatorname{Tr}\left(X^{-1}\right)$, we just consider the Secant linesearch since it is numerically better.
The instances for each problem are split into increasingly smaller subsets depending on their minimum solve time, i.e. the minimum time any of the solvers took to solve it. The cut-offs are at 0 seconds (all problems), took at least 10 seconds to solve, 100 s, 1000 s and lastly 2000 s. Note that if none of the solvers terminates on any instance of a subset, the corresponding row is omitted from the table.
The relative gap is computed for the instances on which at least one method did not terminate within the time limit. That means it excludes the instances on which all methods terminated. The average number of nodes is taken over all solved instances for that solver. The average time is taken using the geometric mean shifted by 1 second. Also note that this is the average time over all instances in that group, i.e. it includes the time outs.
$\operatorname{Tr}\left(X^{-1}\right)$
Monotic + Adaptive
$\operatorname{Tr}\left(X^{-1}\right)$
Secant
$\log\operatorname{Tr}\left(X^{-1}\right)$
Secant
Type
Corr.
Solved
after
inst.
solved
Time (s)
relative
gap
# nodes
solved
Time (s)
relative
gap
# nodes
solved
Time (s)
relative
gap
# nodes
AO no 0 18 50 % 358.32 7.5254 7775 44 % 395.39 7.5383 2369 50 % 279.53 0.3128 3116
10 15 40 % 996.8 9.0286 11600 33 % 1078.18 9.0442 3712 40 % 733.37 0.3736 4637
100 10 10 % 3554.77 13.5379 50135 0 % 3600.14 13.5612 - 10 % 3259.84 0.5553 21991
AO yes 0 18 44 % 831.95 0.1441 39425 44 % 780.03 0.255 44423 61 % 206.45 0.0262 6508
10 13 23 % 2339.34 0.2399 45260 23 % 2358.33 0.3493 50312 46 % 1051.19 0.0326 11869
100 10 0 % 3600.03 0.4124 - 0 % 3600.08 0.4511 - 30 % 2952.66 0.0395 22650
1000 9 0 % 3600.03 0.4019 - 0 % 3600.08 0.4318 - 22 % 3403.62 0.0427 31714
AF no 0 18 83 % 64.13 0.0196 15006 78 % 88.49 0.0198 12429 83 % 17.22 0.0213 355
10 9 67 % 908.07 0.0294 35124 56 % 1373.82 0.0297 31926 67 % 152.17 0.0329 763
AF yes 0 18 28 % 1330.47 2.1036 31537 28 % 1296.61 2.1909 30838 50 % 270.53 5.1204 7754
10 13 15 % 3037.99 2.8795 63405 15 % 2986.85 2.9945 56098 31 % 1837.4 7.9606 17411
100 12 17 % 2995.32 2.9574 63405 17 % 2940.73 3.0822 56098 25 % 2761.53 8.9546 23174
1000 11 9 % 3399.3 3.2253 87165 9 % 3363.42 3.3615 79157 18 % 3015.56 10.2324 23230